### v1

# Detailed nuclear structure calculations for coherent elastic neutrino-nucleus scattering

Abstract
Any discovery of ‘new physics’ in the neutrino sector first requires a precise prediction of the expected Standard Model cross section. Currently, Coherent Elastic neutrino-Nucleus Scattering (CEvNS) experiments are statistics limited. However, as new and future experiments scale up, it will be necessary to improve the theoretical predictions. Here we review the calculation of the CEvNS cross section in a consistent theory of hadronic currents and compute the relevant nuclear form factors using the nuclear shell model. The uncertainty on the form factors is explored by repeating the calculation for various shell model interactions and with Skyme-Hartree-Fock evaluations of the Weak-charge radii. We then refine the Standard Model predictions for the recent experimental results of the COHERENT experiment. We find that our cross sections are in good agreement with previous predictions, but with significantly smaller uncertainties - by up to a factor of 10. Near-future CEvNS experiments will meaningfully benefit from improved predictions through an increased sensitivity to new-physics signals.
I Introduction
The elastic scattering of neutrinos from nuclei via the weak neutral current is a process that can experience a coherent enhancement at sufficiently small momentum transfer. Dubbed CEvNS, the coherent enhancement of this process boosts the cross section by orders of magnitude, making it much larger than other processes involving neutrinos. This enhancement would suggest that the process should be relatively easy to observe, however the observational signature, a low-energy nuclear recoil, is very difficult to detect. Steady advancement in detector technology enabled CEvNS to finally be observed in 2017.
The first detection of neutrinos via CEvNS was performed by the COHERENT experiment using a small (14.6 kg) cesium-iodide detector [1]. This measurement was quickly followed up by a detection using a liquid argon detector [2]. The next planned measurement will be with an array of germanium detectors (18 kg) [3]. All of these measurements were carried out at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory. There, neutrinos with energies up to $\sim$50 MeV are created from pions decaying at rest within the mercury spallation target. Together, these measurements have observed hundreds of CEvNS events and are broadly consistent with the Standard Model expectation for the process.
CEvNS has many potential applications, both practical and fundamental. While the practical applications, such as nuclear reactor monitoring [4], require further detector development before they can be realized, advances in nuclear [5, 6, 7] and fundamental physics are already being realized with current experiments [8, 9, 10, 11]. Notably, CEvNS experiments are the leading probe of some realizations of sterile neutrinos and non-standard neutrino interactions [12]. CEvNS could be detected in dark matter experiments and potentially provide the first warning of a supernova signal [13]. Key to maximising the effectiveness of these applications is having precise predictions for the scattering rate. That is, if one is interested in detecting deviations from the Standard Model at the sub-percent level, then a commensurate level of precision is required for the theoretical predictions.
[…]
In this work we calculate the CEvNS cross sections for nuclear targets used by COHERENT, including the nuclear structure effects, which are typically ignored. To this end, we use large-scale nuclear shell model calculations to derive all relevant nuclear form factors. Using several different nuclear Hamiltonians, we evaluate the uncertainty in the cross section from the nuclear structure effects. These results are then applied to improve the theoretical predictions of the CEvNS event rates that COHERENT has observed to date.
Figure 1: Theoretical predictions of SNS-flux averaged cross sections for the three targets considered in this work. Shown are our main results using the shell model and Skyrme-Hartree-Fock (SHF) and comparisons with the Klein-Nystrand form factor (using electric charge radii) and COHERENT’s prediction. Also shown is COHERENT’s experimental result (black dashed, with gray region showing uncertainty), and a potential future measurement with 5% error (purple hatched).
II CEvNS cross section
Shortly after the discovery of Weak neutral currents in 1973, it was suggested that this interaction should coherently enhance the scattering of neutrinos from the nucleus [19]. At very low momentum transfer, $q$, where $q$ is less than the inverse size of the nucleus, the cross section approximately scales with the number of neutrons in the nucleus. For neutrinos with energies of tens of MeV, the momentum transfer can be large enough to probe the internal structure of the nucleus, resulting in a loss of coherence.
The CEvNS process can be represented by the reaction:
$\nu(p)+\mathcal{N}(k)\rightarrow\nu(p^{\prime})+\mathcal{N}(k^{\prime})$ (1)
where $p,k$ are the incoming momenta of the neutrino and nucleus, respectively, and $p^{\prime},k^{\prime}$ are the outgoing counterparts. The momentum transfer is then given by $t=(p-p^{\prime})^{2}=(k^{\prime}-k)^{2}$. When evaluated in the lab frame (with the nucleus beginning at rest), this gives $t=-q^{2}=-2m_{T}E_{R}$, where $m_{T}$ and $E_{R}$ are the mass and kinetic energy of the recoiling nucleus. With the mass of the neutrinos safely neglected, the maximum recoil an incoming neutrino of energy $E_{\nu}$ can impart is then
$E_{\nu}^{\rm{max}}=\frac{2E_{\nu}^{2}}{m_{T}+2E_{\nu}}.$ (2)
The cross section is typically calculated by treating the nucleus as composed of point-like nucleons and accounting for their spatial extent within the nucleus via a form factor. This results in a standard form of the CEvNS cross section
$\frac{d\sigma}{dE_{R}}=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{E_{R}m_{T}}{2E_{\nu}^{2}}\right)Q_{\text{w}}^{2}F^{2}(q^{2}),$ (3)
where $Q_{\text{w}}=Q_{p}Z+Q_{n}(A-Z)$ is the weak nuclear charge in terms of the proton and neutron weak charges and $F(q)$ is the nuclear form factor, typically taken to be of the Helm [20] or Klein-Nystrand [21] form.
The above is a good approximation to the CEvNS cross section, however, several improvements are variously made in the literature, which can have percent-level corrections. For example, in the above equation, the sub-leading kinematic factor $E_{R}/E_{\nu}$, has been dropped. Additionally, one can take into account the different proton and neutron distributions within the nucleus, assigning them separate form factors [22]. Furthermore, while radiative corrections are typically included in the nucleon weak charges, there are additional flavor-dependent radiative corrections that can be included [23]. Lastly, some effort has been made to include the spin-dependent effects of the neutrinos coupling to odd-$A$ nuclei [24, 25].
Putting these together, but neglecting the spin-dependent effects (which are small but will nevertheless be treated with more rigor in the following section) we can more accurately write the cross section as
$\frac{d\sigma_{T\nu_{i}}}{dE_{R}}=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{E_{R}m_{T}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)$
×$\times\left(Q_{p,\nu_{i}}F_{p}(q^{2})+Q_{n}F_{n}(q^{2})\right)^{2}.$ (4)
Here the flavor dependent charges are taken from [23]:
$Q_{n}$ $=$ $-1.02352(25)$
$Q_{p,\nu_{e}}$ $=$ $0.0747(34),\,\,\,\,Q_{p,\nu_{\mu}}=0.0582(34).$ (5)
Using this equation requires knowledge of the proton and neutron distributions within the nucleus. While the former can be measured well due to its electric charge, the latter is not well constrained experimentally. Given the much larger weak charge of the neutron, the relatively unknown neutron form factor thus contributes the largest uncertainty to determinations of the CEvNS cross section.
III Nuclear form factors
Nuclear form factors are defined as the Fourier transform of density distributions of the nucleus:
$F(q^{2})=\int e^{i\vec{q}.\vec{r}}\rho(r)\ d^{3}\vec{r},$ (6)
where $\rho(r)$ is the relevant density distribution, which is commonly assumed to be a spherically symmetric function. The root-mean-square (rms) radius for a density distribution is then given by
$r^{2}_{\text{rms}}\equiv\langle r^{2}\rangle=\frac{\int\rho(r)\ r^{2}\ d^{3}\vec{r}}{\int\rho(r)\ d^{3}\vec{r}}.$ (7)
The density distribution used, i.e. electric charge-, weak charge- or mass-density, results in different form factors (and rms radii) which are appropriate for use in different scattering processes. Neutral-current neutrino scattering proceeds through the coupling of the Z-boson to the weak charge of the nucleus. Therefore CEvNS requires the use of weak-charge form factors which in turn require knowledge of the weak-charge density. We note that some past work has used experimentally determined electric-charge rms radii to construct form factors for CEvNS, but this can differ significantly from the weak-charge radii [26].
It is possible to measure the weak charge distribution through parity-violating electron scattering, which has been performed at Jefferson lab by the PREX and CREX experiments, on lead-208 and calcium-48 respectively. However, these experiments have only measured the weak charge at a single momentum transfer: $q=0.76$ fm-1 and 1.28 fm-1, respectively [27, 28]. Additionally, these experiments are difficult and unlikely to be performed on every isotope relevant to CEvNS. Therefore, it is necessary to use indirect knowledge of the weak-charge distribution from e.g. the electric charge distributions, or theoretical modelling.
In this work we will compare several methods for modelling the CEvNS form factors and determine their effect on the resulting cross sections. This includes theoretical modelling using both Skyrme-Hartree-Fock (SHF) calculations and the nuclear shell model. The SHF results will allow us to accurately model the nucleon distributions, from which rms radii are computed. The shell model calculations will allow us to include sub-leading nuclear effects through the chiral EFT formalism of [15]. For comparison with previous work in the literature, we will also perform a data-driven calculation based on electric-charge radii.
When modelling the nucleus as a bound state of point-like nucleons, the neutron and proton point radius distributions $R_{n},p$ are obtained from
$R_{n}^{2}=\frac{1}{A-Z}\int d^{3}r\,r^{2}\rho_{n}(\mathbf{r}),$ (8)
$R_{p}^{2}=\frac{1}{Z}\int d^{3}r\,r^{2}\rho_{p}(\mathbf{r}),$ (9)
respectively. Given that stable nuclei exhibit a common, almost uniform, saturating density it is possible to parameterise their density distributions (to very good accuracy) with fits to distributions of just two parameters: one which defines the size of the nucleus and another which defines the thickness of the ‘skin’ region at the surface. At low momentum transfers, the form factor remains relatively insensitive to differences in these parameterizations. To fit observation to a density distribution, it can be sufficient to equate a measured or theoretically determined rms radius with that of the parameterized distribution. We will make use of this matching process below.
III.1 From electric-charge densities
The most reliable access we have to the nuclear size and shape is via electromagnetic probes which couple to the electric charge distribution. Protons, being charged, contribute more to this distribution than the neutrons do. However, the reverse is true for the weak-charge distribution where neutrons dominate in their contribution. This makes the electric charge distribution an imperfect surrogate for the weak-charge distribution. To account for this difference, previous works have made some assumptions about the neutron distribution based on the proton distribution. In [22] the neutron rms radius was chosen to lie between the proton rms charge radius and an additional 0.3 fm. In [29], a fit is made to the proton-neutron radii difference data, however the large scatter in the data used in the fit tends to exaggerate the uncertainty for any one nucleus.
In this work we will follow the method of [22]. For our electric-charge inspired distributions we will use the Klein-Nystrand (KN) [21] model which has been previously used in CEvNS research (including by the COHERENT collaboration). The Klein-Nystrand form factor is based on both a Yukawa potential of range $a_{k}$ and a hard sphere distribution with radius $r_{A}$. It has the form
$F_{\text{KN}}(q)=3\frac{\sin(qr_{A})-qr_{A}\cos(qr_{A})}{(qr_{A})^{3}}\left[1+(qa_{k})^{2}\right]^{-1},$ (10)
where $a_{k}=0.7$ fm and $r_{A}$ is the nuclear radius. The nuclear radius is obtained by equating the rms radius radius for the distribution:
$\langle r^{2}\rangle_{KN}=\frac{3}{5}r_{A}^{2}+6a_{k}^{2}.$ (11)
with the proton rms charge radius. The proton rms charge radii for the target nuclei $\sqrt{\langle r_{p}^{2}\rangle}\equiv r_{\text{rms}}^{p}$ can be obtained from various experimental sources. The values we adopt are shown in Tab. 1. The neutron rms radii are then taken to lie in the range $r_{\text{rms}}^{p}\leq r_{\text{rms}}^{n}\leq r_{\text{rms}}^{p}+0.3$ fm for Ge, I and Cs, whilst for Ar the upper limit is taken to be $r_{\text{rms}}^{p}+0.2$ fm. We use this range to bracket the uncertainty in the neutron distribution.
Table 1: Experimental charge rms radii $r_{\text{rms}}^{p}$ values (in fm) for 40Ar, 70,72,73,74,76Ge, 127I and 133Cs, obtained from [30].
40Ar Ge 127I 133Cs
70 72 73 74 76
$r_{\text{rms}}^{p}$ 3.4274(26) 4.0414(12) 4.0576(13) 4.0632(14) 4.0742(12) 4.0811(12) 4.7500(81) 4.8041(46)
III.2 From Skyrme-Hartree-Fock distributions
To model the weak-charge distribution of the nucleus more accurately, we perform Skyrme-Hartree-Fock calculations to find the point-like nucleon distributions within the nucleus. While we have experimental probes of the electric-charge distributions, which dominantly probes the protons, for consistency we will also use the theoretical values for $R_{p}$ in this calculation. The weak radius of the nucleus can then be calculated from:
$R_{W}^{2}=\frac{1}{Q_{\text{w}}}\left(ZQ_{p}(R_{p}^{2}+r_{\text{w},p}^{2})+(A-Z)Q_{n}(R_{n}^{2}+r_{\text{w},n}^{2})\right)$ (12)
where $r_{\text{w},p}=1.545(17)$ fm and $r_{\text{w},n}=0.8434(12)$ fm are the weak radii of the proton and neutron, respectively.
The nucleon densities were evaluated with the skyax solver [31] assuming axial symmetry for several parametrizations of the Skyrme functional, namely SkM, SLy6, SV-min and UNEDF1 as defined in skyax. Pairing correlations are treated at the BCS level with a “density-dependent delta interaction” (DDDI) for SV-min and SkM, and with a “volume delta interaction” (VDI) for SLy6 and UNEDF1. The final theoretical values of $R_{p,n}$ are obtained from the average of the results for each interaction, while the standard deviations are used to evaluate the uncertainties. The full results for each interaction are compiled in the appendix (Table 6). From this we find the average and standard deviation for each isotope and use it to compute weak radii from Eq. (12). The results are summarised in Table 2.
For simplicity we then parameterize the distribution via the symmetrized 2-parameter Fermi (Fermi) model [32], which has been shown to provide a reasonable fit to similar calculations [26]. The Fermi distribution is defined by:
$\rho(r)=\rho_{0}\frac{\sinh(c/a)}{\cosh(r/a)+\cosh(c/a)},$ (13)
where its two parameters $c$ and $a$ relate to the size and thickness of the skin, respectively. The rms radius of this distribution is:
$\langle r^{2}\rangle=\frac{3}{5}c^{2}+\frac{7}{5}\pi^{2}a^{2},$ (14)
and the normalization constant is
$\rho_{0}=\frac{3Q_{\text{w}}}{4\pi c(c^{2}+\pi^{2}a^{2})}.$ (15)
The (normalised) form factor is then:
$F_{F}(q)$ $=\frac{3}{qc}\left(\frac{\sin(qc)}{(qc)^{2}}\frac{\pi qa}{\tanh(\pi qa)}-\frac{\cos(qc)}{qc}\right)$ (16)
×$\times\frac{\pi qa}{\sinh(\pi qa)}\frac{1}{1+(\pi a/c)^{2}}$
Motivated by the results for the weak skin of calcium and lead in [26], in this work we set $a=0.55\pm 0.05$ fm. We then derive the parameter $c$ by equating $R_{W}=\langle r^{2}\rangle$, i.e. Eq. (14) with Eq. (12).
Table 2: Skyrme-Hartree-Fock predictions of rms point-like proton and neutron radii and the resulting predicted RMS Weak radius.
Nucleus $R_{p}$ (fm) $R_{n}$ (fm) $R_{W}$ (fm)
40Ar 3.326(25) 3.420(17) 3.514(17)
70Ge 3.930(21) 3.986(12) 4.065(12)
72Ge 3.946(20) 4.036(12) 4.116(12)
73Ge 3.954(19) 4.059(11) 4.140(12)
74Ge 3.963(18) 4.084(10) 4.165(11)
76Ge 3.993(20) 4.143(14) 4.224(15)
127I 4.688(14) 4.824(9) 4.895(10)
133Cs 4.750(12) 4.887(8) 4.957(9)
IV Detailed nuclear treatment
The formalism discussed thus far has been derived top-down, dealing with the most dominant contributions of neutrino-nucleus couplings. However, this does not systematically account for the nuclear effects stemming from the neutrino coupling to both vector (spin-independent) and axial-vector (spin-dependent) currents within the nucleus. These effects are generally expected to be small and therefore are typically ignored.
The full, bottom-up, calculation starts with the hadronic currents within the nucleus and using effective field theory, derives the Wilson coefficients for the nuclear couplings. This task has been performed for the CEvNS process, in full generality, in [15]. The resulting cross section is much like the traditional form of Eq.(3), with an additional term for the axial form factor:
$\begin{split}\frac{d\sigma_{A}}{dE_{R}}&=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{m_{T}E_{R}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)Q_{\text{w}}^{2}|F_{\text{w}}(\mathbf{q}^{2})|^{2}\\ &+\frac{G_{F}^{2}m_{T}}{4\pi}\left(1+\frac{m_{T}E_{R}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)F_{A}(\mathbf{q}^{2}),\end{split}$ (17)
However, the full form of the nuclear form factors go beyond the density distributions discussed previously. The Weak form factor takes the form
$\begin{split}F_{\text{w}}(\mathbf{q}^{2})&=\frac{1}{Q_{\text{w}}}\bigg{[}\bigg{(}Q_{\text{w}}^{p}\Big{(}1+\frac{\langle r_{E}^{2}\rangle^{p}}{6}t+\frac{1}{8m_{N}^{2}}t\Big{)}\\ &+Q_{\text{w}}^{n}\frac{\langle r_{E}^{2}\rangle^{n}+\langle r_{E,s}^{2}\rangle^{N}}{6}t\bigg{)}\mathcal{F}^{M}_{p}(\mathbf{q}^{2})\\ &+\bigg{(}Q_{\text{w}}^{n}\Big{(}1+\frac{\langle r_{E}^{2}\rangle^{p}+\langle r_{E,s}^{2}\rangle^{N}}{6}t+\frac{1}{8m_{N}^{2}}t\Big{)}\\ &+Q_{\text{w}}^{p}\frac{\langle r_{E}^{2}\rangle^{n}}{6}t\bigg{)}\mathcal{F}^{M}_{n}(\mathbf{q}^{2})\\ &-\frac{Q_{\text{w}}^{p}(1+2\kappa^{p})+2Q_{\text{w}}^{n}(\kappa^{n}+\kappa_{s}^{N})}{4m_{N}^{2}}t\mathcal{F}^{\Phi^{\prime\prime}}_{p}(\mathbf{q}^{2})\\ &-\frac{Q_{\text{w}}^{n}(1+2\kappa^{p}+2\kappa_{s}^{N})+2Q_{\text{w}}^{p}\kappa^{n}}{4m_{N}^{2}}t\mathcal{F}^{\Phi^{\prime\prime}}_{n}(\mathbf{q}^{2})\bigg{]},\end{split}$ (18)
where $m_{N}$ is the nucleon mass and the remaining constants are provided in [15]. The weak form factor hence depends on the coherent and semi-coherent responses $F^{M}_{N}$ and $F^{\Phi^{\prime\prime}}_{N}$, respectively, where $N=\{p,n\}$ denotes protons and neutrons. The axial-vector form factor has the form
$\begin{split}F_{A}&(\mathbf{q}^{2})=\frac{8\pi}{2J+1}\\ \times&\left((g_{A}^{s,N})^{2}S_{00}^{\mathcal{T}}(\mathbf{q}^{2})-g_{A}g_{A}^{s,N}S_{01}^{\mathcal{T}}(\mathbf{q}^{2})+(g_{A})^{2}S_{11}^{\mathcal{T}}(\mathbf{q}^{2})\right)\end{split}$ (19)
where, in the isospin-basis, we have the spin structure functions in terms of the nuclear responses ±$\mathcal{F}_{\pm}^{\Sigma^{\prime}_{L}}=\mathcal{F}_{p}^{\Sigma^{\prime}_{L}}\pm\mathcal{F}_{n}^{\Sigma^{\prime}_{L}}$:
$\begin{split}S_{00}^{\mathcal{T}}&=\sum_{L}\left[\mathcal{F}_{+}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\right]^{2},\\ S_{11}^{\mathcal{T}}&=\sum_{L}\left[[1+\delta^{\prime}(\mathbf{q}^{2})]\mathcal{F}_{-}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\right]^{2},\\ S_{01}^{\mathcal{T}}&=\sum_{L}2\left[1+\delta^{\prime}(\mathbf{q}^{2})\right]\mathcal{F}_{+}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\mathcal{F}_{-}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2}),\end{split}$ (20)
The correction to the leading SD coupling for the transverse SD response $\delta^{\prime}(\mathbf{q}^{2})$ is given in [15]. In this more-detailed treatment of the nucleus, the full Weak radius receives additional contributions from spin-orbit effects and the strange electric-charge distribution. It takes the form
$\begin{split}R_{\text{w}}^{2}&=\frac{ZQ_{\text{w}}^{p}}{Q_{\text{w}}}\bigg{(}R_{p}^{2}+\langle r_{E}^{2}\rangle^{p}+\frac{Q_{\text{w}}^{n}}{Q_{\text{w}}^{p}}\left(\langle r_{E}^{2}\rangle^{n}+\langle r_{E,s}^{2}\rangle^{N}\right)\bigg{)}\\ &+\frac{NQ_{\text{w}}^{n}}{Q_{\text{w}}}\bigg{(}R_{n}^{2}+\langle r_{E}^{2}\rangle^{p}+\langle r_{E,s}^{2}\rangle^{N}+\frac{Q_{\text{w}}^{p}}{Q_{\text{w}}^{n}}\langle r_{E}^{2}\rangle^{n}\bigg{)}\\ &+\frac{3}{4m_{N}^{2}}+\langle\tilde{r}^{2}\rangle_{\text{so}},\end{split}$ (21)
where the spin-orbit contribution is given by:
$\begin{split}\langle\tilde{r}^{2}\rangle_{\text{so}}&=-\frac{3Q_{\text{w}}^{p}}{2m_{N}^{2}Q_{\text{w}}}\Big{(}1+2\kappa^{p}+2\frac{Q_{\text{w}}^{n}}{Q_{\text{w}}^{p}}(\kappa^{n}+\kappa_{s}^{N})\Big{)}\mathcal{F}^{\Phi^{\prime\prime}}_{p}(0)\\ &-\frac{3Q_{\text{w}}^{n}}{2m_{N}^{2}Q_{\text{w}}}\Big{(}1+2\kappa^{p}+2\kappa_{s}^{N}+2\frac{Q_{\text{w}}^{p}}{Q_{\text{w}}^{n}}\kappa^{n}\Big{)}\mathcal{F}^{\Phi^{\prime\prime}}_{n}(0).\end{split}$ (22)
To apply this formalism and calculate cross sections for a given target we will employ the nuclear shell model to compute the nuclear responses above.
IV.1 Evaluating the nuclear responses
Figure 2: Spin-independent weak form factors for CEvNS in the four targets relevant to current experiments. Bands show the error due to uncertainties in the neutron distribution (Charge) and from using different interactions (SHF and Shell model).
To compute the weak and axial-vector nuclear form factors in Eqs. (18) and (19), we need to evaluate the $\mathcal{F}^{M}_{N}$, $\mathcal{F}^{\Phi^{\prime\prime}}_{N}$ and $\mathcal{F}^{\Sigma^{\prime}}_{N}$ nuclear responses. The nuclear responses originate from a multipole decomposition of the lepton-nucleus amplitude, developed in [33, 34]. Here, they are defined as:
$|\mathcal{F}^{X}_{N}(q^{2})|^{2}\equiv$ $\frac{4\pi}{2J_{i}+1}\sum\limits_{J=0}\langle J_{i}||X_{J}^{(N)}||J_{i}\rangle\langle J_{i}||X^{(N)}_{J}||J_{i}\rangle,$
$|\mathcal{F}^{\Sigma^{\prime}}_{N}(q^{2})|^{2}\equiv$ $\sum\limits_{J=1}\langle J_{i}||\Sigma_{J}^{\prime(N)}||J_{i}\rangle\langle J_{i}||\Sigma^{\prime(N)}_{J}||J_{i}\rangle,$
where $J_{i}$ is the nuclear ground state angular momentum, and the nuclear operators $X=M,\,\,\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are given in Appendix C.
The nature of these operators can be understood from their long-wavelength ($q\rightarrow 0$) limit (provided in Table 3). The $M$ operator is spin-independent (SI) and adds coherently from all nucleons, giving rise to the standard coherent enhancement of CEvNS. The $\Phi^{\prime\prime}$ term in the $q\rightarrow 0$ limit is proportional to a spin-orbit operator, and is sensitive to nucleon occupations of the spin-orbit partner levels. This partially coherent operator is highest when one spin-orbit partner is filled and the other is empty. The $\Sigma^{\prime}$ operator is spin-dependent (SD), picking out the transverse projection of the spin current. All nuclei have non-zero responses to the $M$ and $\Phi^{\prime\prime}$ operators, while only odd-nuclei with a $J_{i}\neq 0$ ground state have a $\Sigma^{\prime}$ response.
Table 3: Leading order terms of the nuclear operators in the long-wavelength limit $q\rightarrow 0$ [35].
Response Type Leading Multipole Long-wavelength Limit
$M_{JM}$ : Charge $M_{00}(q\vec{x}_{m})$ $\frac{1}{\sqrt{4\pi}}1(m)$
$T^{\text{el5}}_{JM}$ : Axial Transverse Electric $\Sigma^{\prime}_{1M}(q\vec{x}_{m})$ $\frac{1}{\sqrt{6\pi}}\sigma_{1M}(m)$
$L_{JM}$ : Longitudinal $\frac{q}{m_{N}}\Phi^{\prime\prime}_{00}(q\vec{x}_{m})$ $-\frac{q}{3m_{N}\sqrt{4\pi}}\vec{\sigma}(m).\vec{l}(m)$
$\frac{q}{m_{N}}\Phi^{\prime\prime}_{2M}(q\vec{x}_{m})$ $-\frac{q}{m_{N}}\frac{1}{\sqrt{30\pi}}\left[x_{m}\otimes\left(\vec{\sigma}(m)\times\frac{\vec{\nabla}}{i}\right)_{1}\right]_{2M}$
Using the Wigner-Eckhart theorem the nuclear matrix elements in Eq. (IV.1) can be reduced in angular momentum and isospin:
$\langle J_{i};TM_{T}\ ||\sum_{m=1}^{A}\hat{O}_{J,\tau}(q\vec{x}_{m})||\ J_{i};TM_{T}\rangle=(-1)^{T-M_{T}}$
$\begin{pmatrix}\vspace{2mm}T&\tau&T\\ \vspace{2mm}-M_{T}&0&M_{T}\hskip 5.69054pt\end{pmatrix}\langle J_{i};T\ \vdots\vdots\sum_{m=1}^{A}\hat{O}_{J,\tau}(q\vec{x}_{m})\vdots\vdots\ J_{i};T\rangle.$
Here, $\hat{O}_{J,\tau}$ signify the nuclear operators, with $\hat{O}_{J,\tau}=\hat{O}_{J}\ \tau_{3}^{\tau}$ and $\tau_{3}$ being the nucleon isospin operator ($\tau=\{0,1\}$). The isospin-independent(dependent) component is given by the $\tau=0$ ($\tau=1$) term.
This can then be written as a product of single-nucleon matrix elements and one-body density matrix elements (OBDMEs) $\Psi^{J;\tau}_{|\alpha|,|\beta|}$ in the harmonic oscillator basis [36]:
$\langle J_{i};TM_{T}\ ||\sum_{m=1}^{A}\hat{O}_{J,\tau}(q\vec{x}_{m})||\ J_{i};TM_{T}\rangle=(-1)^{T-M_{T}}$
×$\times\begin{pmatrix}\vspace{2mm}T&\tau&T\\ \vspace{2mm}-M_{T}&0&M_{T}\hskip 5.69054pt\end{pmatrix}\sum_{|\alpha|,|\beta|}\Psi^{J;\tau}_{|\alpha|,|\beta|}\langle|\alpha|\ \vdots\vdots\hat{O}_{J,\tau}(q\vec{x})\vdots\vdots\ |\beta|\rangle.$
[…]
$\begin{split}\Psi^{J;\tau}_{|\alpha|,|\beta|}&\equiv\frac{\langle J_{i};T\ \vdots\vdots\left[a^{\dagger}_{|\alpha|}\otimes\tilde{a}_{|\beta|}\right]_{J;\tau}\vdots\vdots\ J_{i};T\rangle}{\sqrt{(2J+1)(2\tau+1)}},\end{split}$ (26)
where $\tilde{a}_{\beta}=(-1)^{j_{\beta}-m_{j_{\beta}}+1/2-m_{t_{\beta}}}\ a_{|\beta|;-m_{j_{\beta}},-m_{t_{\beta}}}$ and $\otimes$ denotes a tensor product. This basis gives response functions of the form $e^{-2y}p(y)$, where $p(y)$ is a polynomial with $y=(qb/2)^{2}$ and $b={1}/{\sqrt{m_{N}\omega}}\approx\sqrt{41.467/(45A^{-1/3}-25A^{-2/3})}$ fm is the harmonic oscillator length parameter ($\omega$ is the oscillator frequency).
IV.2 Form factors from shell model calculations
Large-scale nuclear shell model calculations were performed using NuShellX [37], which is often employed to obtain theoretical energies, wave functions and spectroscopic overlaps for a wide range of nuclei.The program’s inputs include the relevant valence space for the nucleus/isotope of interest - which constitutes a set of single nucleon levels above a completely filled nuclear core - as well as its corresponding list of two-body interactions. Each interaction is fitted to a set of experimental nuclear data for a range of nuclei in the relevant valence space. The contemporary approach is to derive the interactions from the nucleon-nucleon interaction using effective field theory. The interactions are then fine tuned by fitting a limited number of parameters to a body of nuclear data - usually excitation energies in a range of nuclei in the valence space. The aforementioned OBDMEs are extracted for the valence space levels, with the values dependent on the particular shell model interactions used.
In this work we explore a variety of shell model interactions and truncations for our chosen targets: 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs. Using the shell model calculations detailed below, we extract the OBDME for all targets. The OBDMEs are then used to compute the nuclear responses using the Mathematica package described in [38]. Analytic form factor expressions for 40Ar, 70,72,73,74,76Ge and 127I are provided in [39], while those for 133Cs are provided in Appendix D. The form factors for each of the nuclear responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are plotted as functions of $q$ in Appendix D. The form factors are used to derive point-proton, point-neutron and weak radii for all isotopes for comparison with those obtained with our SHF calculations, and are given in Table 7 of Appendix A. Additionally, we make comparisons with experimental energy spectra, moments and transitions in Appendix B.
The shell-model derived Weak form factors are shown in Fig. 2, compared to form factors based on the charge distribution and SHF calculations. The uncertainty band denotes the range of values obtained given the different shell model interactions. The shell model results for each isotope show good agreement with each other (low overall uncertainty) and fall within the large uncertainty band of the form factors derived from the charge distribution. They are also generally in very good agreement with the SHF results within the experimentally accessible regions of interest. The axial form factor is shown in Fig. 3 for each interaction individually.
The sections below outline the valence space, valence space truncation and shell model interactions used in this work for each target.
Figure 3: Spin-dependent form factors relevant to CEvNS for the three $J_{i}\neq 0$ targets we consider. The curves denote the different shell model interactions as described in the text.
40Ar
The $sdpf$ valence space was employed for this nucleus, which constitutes single particle levels $1d_{5/2}$, $2s_{1/2}$, $1d_{3/2}$, $1f_{7/2}$, $2p_{3/2}$, $1f_{5/2}$ and $2p_{1/2}$. The valence space truncation employed consisted of an unrestricted $sd$ shell and an empty $pf$ shell for the protons, and a full $sd$ shell with an unrestricted $pf$ shell for the neutrons. Four shell model interactions were used, namely SDPF-NR [40], SDPF-U [41], EPQQM [42] and SDPF-MU [43], all of which consider a 16O core. This isotope was not considered in the work of [35, 38].
The 40Ar theoretical energy levels for the above shell model interactions are plotted against the experimental values in Fig. 5. The first two theoretical energy levels for all interactions are consistent with experiment within $\sim 200$ keV, except for SDPF-MU which is within $\sim 400$ keV from experiment for the $2_{1}^{+}$ state. Overall, the energy spectrum of the EPQQM interaction is least consistent with the other shell model calculations and with experiment, with the spectrum being much more sparse and spread out in the level spacings.
The magnetic dipole and electric quadrupole moments and transitions are presented in Table 8, where the effective $g$ factors and charges employed take on values of $e_{p}=1.5$, $e_{n}=0.5$, $g_{sp}=5.586$, $g_{sn}=-3.826$, $g_{lp}=1$, and $g_{ln}=0$. Information on the ground state is most important for comparison with theoretical values, as it is used in the scattering calculations here. However, as the ground state of 40Ar is $0^{+}$, the magnetic dipole and electric quadrupole moment values are identically zero. Nevertheless, we employ the next non-zero state ($2_{1}^{+}$), which can give an indication of the overall accuracy of the calculation through information on the spectrum beyond the ground state. It is found that the shell model moments are not consistent with experimental values, and most experimental transitions are poorly reproduced as well.
The reason for the failure of the shell model for the case of 40Ar is understood: the assumption that 40Ca is an inert core is not justified. Its ground state is near spherical, as assumed in the shell model, but it has low-excitation deformed states formed by exciting pairs of protons and neutrons across the shell gaps at $N=20$ and $Z=20$. This leads to mixing between the spherical shell model states and the deformed states in nuclei like 40Ar. The case of 42Ca, an isotone of 40Ar, is best studied, as has recently been reviewed by Stuchbery and Wood [44]. They infer that the first 2+ state of 42Ca is about 50% due to the spherical shell model configuration. The ground state is about 65% of the shell model configuration. Similar mixing is present in 40Ar. Thus the shell model calculations performed here are limited by the presence of configuration mixing well beyond the usual valence space.
The obtained form factors are compared against those based on the SHF calculations and the charge distribution in Appendix D. The maximum difference between the form factors for the $M$ response at the minimum $q$ value is $\sim 0.3\%$, and $\sim 11\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 24\%$ and $\sim 21\%$, respectively, where the largest differences occur between the SDPF-MU and EPQQM interactions.
70,72,73,74,76Ge
The germanium shell model calculations were performed in the $f_{5}pg_{9}$ model space, which consists of the single-particle levels $2p_{3/2}$, $1f_{5/2}$, $2p_{1/2}$ and $1g_{9/2}$. The work of [35, 38] employed the GCN2850 interaction [45], with a valence space truncation that consisted of limiting the occupation number of the $1g_{9/2}$ level to no more than two nucleons above the minimum occupation for all isotopes. Here, we consider an unrestricted $f_{5}pg_{9}$ model space with the JUN45 [46] and jj44b [47] interactions.
The experimental energy spectra are plotted against the theoretical shell model ones for the even isotopes in Fig. 6 (Appendix B.2). The spectra comparison for 73Ge is provided in [39], alongside a brief summary of the two aforementioned shell model interactions, where the theoretical spectra were found to be inconsistent with another, and not particularly consistent with experiment. In the cases of 70,72,74Ge, the low-lying experimental spectra are a little better reproduced by the JUN45 interaction results, which was fitted to the Ge isotopes in consideration (the jj44b interaction on the other hand was not fitted to these). The shell model spectra for 76Ge are practically identical to one another up to $\sim 1.9$ MeV. An agreement between theoretical and experimental levels of $\sim 200$ keV would be considered satisfactory for such shell model calculations using heavier nuclei, however this agreement exists only for some levels of the even isotopes.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided for the even isotopes in Tables 9 and 10. Much like with 40Ar, the magnetic and electric moment information is not available for the ground state $0^{+}$, however the subsequent non-zero states are used here as a test of the overall accuracy of the shell model code. As before, the values for 73Ge are provided in [39]. For the magnetic dipole moments of the even isotopes, the experimental $2_{1}^{+}$ and $4_{1}^{+}$ states are well reproduced by both shell model interactions, whereas the jj44b value of the $2_{2}^{+}$ state is consistently closer to experiment compared to JUN45. The $2_{1}^{+}$ electric quadrupole moment is quite inconsistent with the experimental values for 70,72,74Ge, where only the jj44b 76Ge value lies within the experimental range.
The individual germanium form factors are plotted (for the $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ responses) in Appendix D, and compared against the SHF and charge distribution form factors. The $M$ response is consistent amongst all calculations, with a maximum difference of $\sim 0.2\%$ at the lower end of the COHERENT $q$ region-of-interest (ROI) for the even isotopes, and $\sim 3\%$ at the higher end for 70,72Ge. For 73Ge these take on the values of $\sim 0.3\%$, and $\sim 20\%$ at the higher end for 73,74,76Ge. For the $\Phi^{\prime\prime}$ channel, the maximum difference is observed between the GCN2850 and jj44b shell model calculations, which is $\sim 80\%$ for 70Ge at both ends of the $q$ spectrum, factor of $\sim 2.4$ for 72Ge, factor of $\sim 2.3$ for 73Ge, factor of $\sim 2.2$ for 74Ge, and $\sim 60\%$ for 76Ge. The difference in the SD $\Sigma^{\prime}$ channel for 73Ge is $\sim 30\%$ at both ends of the $q$ ROI.
127I
The iodine calculations were performed in the model space which includes all proton and neutron orbits in the major shell between magic numbers 50 and 82, with single particle levels $1g_{7/2}$, $2d_{5/2}$, $2d_{3/2}$, $3s_{1/2}$ and $1h_{11/2}$. The work of [35, 38] utilises a nuclear shell model interaction developed by Baldridge and Dalton [48], which we refer to as “B&D” [39]. The valence space restriction employed in [35, 38] involves fixing the occupation number of $1h_{11/2}$ to the minimum allowed nucleon number.
Here, we consider the SN100PN [49] and GCN5082 [45] interactions, and perform two calculations for each. Our shell model calculations are performed in the same model space as the B&D calculations of [35, 38], but with different restrictions. The model space restrictions employed in our work are the same as those in [39], with the four different calculations given the names “SN100PN” and “GCN5082” (which share the same truncation), as well as “Truncated SN100PN” (with a more truncated scheme), and “GCN5082 Expanded” (with an alternative truncation to “GCN5082”). The theoretical energy spectra are plotted against the experimental ones in [39], in addition to the theoretical and experimental electric quadrupole and magnetic dipole moments, with the proton and neutron effective charges and $g$ factors also provided in [39].
The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and charge distribution form factors in Appendix D. The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 30\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 55\%$ and $\sim 50\%$, respectively. For $\Sigma^{\prime}$, the factor differences are $\sim 44$ and $\sim 81$, due to the large difference between the Truncated SN100PN calculation and the remaining calculations. This is consistent with the integrated form factor (IFF) results in [39], which showed a more pronounced difference between the SD Truncated SN100PN IFF results and other calculations. This indicates that a more severe valence space truncation for a shell model interaction has a larger effect compared to a second more comparable truncation.
133Cs
Here, we consider the same valence space and shell model interactions (SN100PN and GCN5082) as 127I above. We have performed four calculations for 133Cs, two of which employ the same truncations as the “SN100PN” and “GCN5082” calculations in Section IV (and are given the same names here). We consider two additional calculations, one each for the SN100PN and GCN5082 interactions, where the neutron $2d_{5/2}$ level now has a minimum of 4 neutrons (as opposed to previously being completely filled), with the remainder of the proton and neutrons levels retaining the previous truncations. This is to explore the effect of a less truncated valence space on the final results. These two additional calculations are termed “SN100PN Expanded” and “GCN5082 Expanded”.
The experimental energy spectra are plotted against the theoretical values for all the above calculations in Fig. 7. Only the SN100PN result reproduces the correct ground state spin and parity of $7/2^{+}$, however all four calculations reproduce the correct spins and parities of the three lowest-lying states within $200$ keV of experiment. For both shell model interactions, using a less truncated valence space causes the spacing of the levels to widen compared to the original calculations, generally pushing the spectrum further up. This is likely due to the opening up of the $2d_{5/2}$ level in these calculations.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided in Tables 11 and 12. The proton and neutron effective charges and $g$ factors employed here are the same as those used in [39] for 127I. The experimental magnetic dipole moments are much better reproduced by the shell model calculations compared to the electric quadrupole ones. The experimental electric quadrupole transitions are only reproduced within the uncertainty margins by the GCN5082 calculation, specifically for transitions from the $5/2_{1}^{+}$ and $3/2_{1}^{+}$ states to the ground state. On the other hand, the theoretical magnetic dipole transition values are not consistent with experiment. However, the important test of the shell model wavefunction is the magnetic dipole moment, which is sensitive to the wavefunction of the single state, and to the single nucleon structure of the state, rather than to emerging collective properties, which affect the E2 transition rates and the quadrupole moments. The magnetic dipole transitions are suppressed among the dominant low-excitation configurations, as is evident from their small values, and so become very sensitive to small component in the wavefunction that are not of interest here. Despite the lack of agreement for some observables, the present shell model calculations can be trusted to include the important single-particle configurations in the ground-state wavefunction.
The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and KN ones in Appendix D. The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 20\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 6\%$ and $\sim 7\%$, respectively. For $\Sigma^{\prime}$, these differences are $\sim 13\%$ and $\sim 18\%$.
Table 4: Our CEvNS cross section predictions in units of 10-40cm2
COHERENT
Target Charge SHF Shell Mod. Prediction Exp.
Ar 18.65(22) 18.68(6) 18.64(5) 18.0(4) 23(7)
Ge 59(1) 59.8(2) 59.2(2) - -
CsI 189(4) 189.3(5) 186.3(5) 189(6) 165(30)
V Experimental observables
To date, CEvNS has been observed by the COHERENT collaboration using the stopped pion source of neutrinos from the SNS, with both cesium-iodide [50] and argon [2] detector targets. Presently COHERENT have also deployed a germanium detector and measurements are ongoing. There are also experiments aiming to measure CEvNS from reactor neutrinos, including CONUS [51], RICOCHET [52] and NUCLEUS [53]. However, the momentum transfer from reactor neutrinos is very small and thus the form factor has little effect on the cross section. We therefore choose to focus our present analysis on the cesium-iodide, argon and germanium targets of COHERENT and leave reactor based experiments to a future work.
V.1 Flux averaged cross sections
First we calculate the flux averaged cross sections, by performing a weighted average of the cross sections over the three incoming neutrino flavor spectra and the isotopes in the target:
[…]
Here the neutrino spectra, $\frac{d\phi}{dE_{\nu}}$, are normalized to unity. For this calculation we use the neutrino spectra provided in [54], which were derived from detailed GEANT4 simulations of the proton beam colliding in the mercury spallation target. The uncertainties in the cross sections are determined from two sources: the form factors and the nucleon charges. We neglect uncertainties in the modelling of the neutrino spectra, which are unavailable, but note that errors in the spectra should affect all of the compared cross sections approximately the same. The form factor uncertainties are found by computing the flux averaged cross section with our two extremal form factors for each model (i.e. the upper and lower limits of the bands shown in Fig. 2). The central value and uncertainty is then taken as ($\sigma_{\rm{max}}+\sigma_{\rm{min}}$)/2 and ($\sigma_{\rm{max}}-\sigma_{\rm{min}}$)/2, respectively. The uncertainty in the nucleon charges is taken from [23]. The total uncertainty is determined by summing these two sources in quadrature. The KN and SHF models are calculated using Eq. (II), while the shell model result uses the full cross section in Eq. (3). The results of this calculation are shown in Fig. 1 and provided numerically in table 4.
Our argon results are all in very close agreement with each other, owing to the fact that argon is an even-even nucleus which doesn’t receive large spin-orbit, and no spin-dependent, contributions. This result is slightly larger than the COHERENT prediction, though with the large uncertainties, is still in reasonable agreement. Our detailed nuclear results do shrink the uncertainty band, though it will be some time before the statistical errors in the experimental results make this difference relevant.
Our cesium-iodide results are in good agreement with the COHERENT prediction. Here we do see a small difference between our shell model results and the SHF calculations. While both cesium and iodine are odd nuclei, the axial term is a positive contribution and so the resulting difference stems from the spin-orbit term.
Our germanium results show a similar trend to the cesium-iodide results, with the shell model results being slightly smaller than the SHF results, which are both with the prior uncertainty derived from charge distributions. Presently there are no COHERENT results to compare with, though they are expected in the near future.
Figure 4: Total event count predictions for the COHERENT experiment’s argon and cesium-iodide detectors, using our 3 cross section calculations. Also shown is COHERENT’s theoretical prediction and experimental result.
[…]
We note that the assumed cross section model will affect the experimentally determined flux averaged cross section through the calculation of the experiment’s sensitivity. Thus as measurements become better, one must be cautious to not interpret deviations of experimental results from predictions which used a different cross section model. For a direct comparison with the observable experimental results we also compute the predicted number of events for COHERENT’s detectors.
V.2 Experimental event counts
The differential scattering rate of neutrinos from nuclei, per target and incoming neutrino, is found by integrating over the incoming neutrino spectra and summing them:
[…]
Here, the neutrino flux is integrated from the minimum neutrino energy that can induce a recoil of energy $E_{R}$:
$E_{\nu}^{\mathrm{min}}=\sqrt{\frac{E_{R}m_{T}}{2}}.$ (29)
To go from the raw scattering rate to the expected number of counts in an experiment we follow the recipes provided by the COHERENT collaboration in [54] and [55]. The total number of detected CEvNS events is found by integrating over the detected energy taking into account the quenching of nuclear recoils, detector resolution effects and the energy dependent detection efficiency. For the cesium-iodide detector this has the form
[…]
×$\times\int dE_{R}P(P\!E|E_{R})\frac{dR}{dE_{\nu}},$
while for the argon detector the form is
[…]
×$\times\int dE_{R}\,P(E_{ee}|E_{R})\frac{dR}{dE_{\nu}},$
where $P\!E$ is the detected photoelectrons, $N_{T}$ is the number of targets in the detector and $\Phi=\frac{\rm{POT}f_{\pi}}{4\pi d^{2}}$ is total number of neutrinos (per flavor) expected at the target [calculated from the total number of protons on target (POT) and the average number of pions produced per POT, $f_{\pi}$]. The values for the experimental parameters are summarized in table 5.
Table 5: Parameters for the COHERENT experimental runs that we considered in this work
Target Mass Distance POT $\langle\pi/\rm{POT}\rangle$
argon 24.4 kg 27.5 m $1.38\times 10^{23}$ 0.09
cesium-iodide 14.6 kg 19.3 m $3.20\times 10^{23}$ 0.0848
The resulting expected counts for the two detectors’ runs are shown in Fig. 4. The uncertainty in the number of counts is determined from two sources: the uncertainty in the cross section (calculated as above), and from the experimental systematic uncertainty (which is dominated by the 10% uncertainty in the total neutrino flux). We take our systematic uncertainty from COHERENT’s analysis (not including their theoretical uncertainties): 13% for argon and 12% for cesium iodide.
Overall we achieved good agreement reproducing COHERENT’s results, with all predictions well within the errors for both COHERENT’s predictions and experimental results. However, without knowledge of their exact theoretical model it is difficult to reproduce the result precisely. In this instance the result obtained using the form factor derived from the electric-charge distribution provides a useful benchmark to compare against. For the cesium-iodide detector our predictions were systematically larger, but shell model results were comparatively lower than from the electric-charge distribution form factors. This is likely due to the slightly larger charge radius predicted from the shell model, which causes coherence to be lost to a greater extent (for a given momentum transfer). For the argon detector our result is systematically lower than COHERENT’s prediction, in contrast to our increased cross section prediction. As with the argon cross sections, the different form factors all produce very similar results because argon-40 is an even-even nucleus.
It is worth noting that the reduction in the theoretical uncertainty achieved for our cross section predictions has essentially no effect on the total uncertainty of our event rate predictions, since they are dominated by the uncertainty in the neutrino flux.
VI Conclusion
In this work we have computed precise flux-averaged cross section predictions for the CEvNS experiments being carried out by the COHERENT collaboration using argon, germanium and cesium-iodide targets. These predictions were made using both Skyme-Hartree-Fock calculations of weak radii, and shell model calculations of nuclear structure factors. The shell model results included sub-leading nuclear effects such as from spin-orbit coupling and axial currents. We used the Skyme-Hartree-Fock calculation to provide a cross-check of the density distribution obtained with the shell model. The results for the $F_{M}$ response showed very good agreement between these two methods.
Our results are in broad agreement with prior work but have substantially reduced the theoretical errors by up to a factor of 10. Future advances on the experimental side will both reduce the uncertainty on the neutrino flux and increase the number of detected CEvNS events, reducing statistical errors. Combined with these future advancements, our results will pave the way for CEvNS to serve as a precision test of the Standard Model.
Acknowledgements
We are grateful to Daniel Pershey, Jacob Zettlemoyer and Oleksandr Tomalak for useful discussions. This work was supported by the Australian Research Council through Discovery Project DP220101727. The work of JLN is supported by the Australian Research Council through the ARC Centre of Excellence for Dark Matter Particle Physics, CE200100008.
Appendix A Point-like nucleon distributions
Full results for the point-like distribution of nucleons in each nuclear target are given in the tables below. The SHF results were used to construct Fermi-form factors for the Weak-charge distribution. The shell model results are given for comparison only and are not used for any subsequent calculations. The shell model point-nucleon radii were calculated using
$R_{p}^{2}=-\frac{6}{Z}\frac{\text{d}F^{M}_{p}(\mathbf{q}^{2})}{\text{d}\mathbf{q}^{2}}\bigg{|}_{\mathbf{q}^{2}=0},$ (32)
with an equivalent form for the neutrons.
Table 6: Skyrme-Hartree-Fock predictions of root mean square point like proton and neutron radii.
Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm)
40Ar SkM 3.356 3.445
SLy6 3.330 3.412
SV-min 3.321 3.418
UNEDF1 3.296 3.407
Average 3.326(25) 3.420(17)
70Ge SkM 3.949 4.003
SLy6 3.944 3.988
SV-min 3.922 3.980
UNEDF1 3.904 3.975
Average 3.930(21) 3.986(12)
72Ge SkM 3.965 4.053
SLy6 3.959 4.031
SV-min 3.940 4.031
UNEDF1 3.922 4.028
Average 3.946(20) 4.036(12)
73Ge SkM 3.973 4.077
SLy6 3.965 4.054
SV-min 3.947 4.054
UNEDF1 3.931 4.053
Average 3.954(19) 4.059(11)
74Ge SkM 3.980 4.099
SLy6 3.976 4.081
SV-min 3.955 4.077
UNEDF1 3.941 4.078
Average 3.963(18) 4.084(10)
76Ge SkM 4.016 4.163
SLy6 4.004 4.138
SV-min 3.978 4.129
UNEDF1 3.975 4.140
Average 3.993(20) 4.143(14)
127I SkM 4.702 4.836
SLy6 4.698 4.824
SV-min 4.676 4.814
UNEDF1 4.676 4.822
Average 4.688(14) 4.824(9)
133Cs SkM 4.763 4.898
SLy6 4.759 4.885
SV-min 4.739 4.878
UNEDF1 4.740 4.887
Average 4.750(12) 4.887(8)
Table 7: Shell model predictions of the point-proton, point-neutron and weak radii, using the various shell model calculations for each of the isotopes considered. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm) $R_{\text{w}}$ (fm)
40Ar SDPF-NR 3.32852 3.43528 3.53293
SDPF-U 3.32852 3.43528 3.53282
SDPF-MU 3.32851 3.43528 3.53243
EPQQM 3.32851 3.43528 3.53348
70Ge GCN2850 3.95524 4.05401 4.13886
JUN45 3.95843 4.05818 4.14331
jj44b 3.95963 4.07061 4.15307
72Ge GCN2850 3.97233 4.07882 4.16942
JUN45 3.97412 4.11049 4.19586
jj44b 3.97457 4.12156 4.20457
73Ge GCN2850 3.97979 4.11144 4.20019
JUN45 3.98047 4.13135 4.21644
jj44b 3.97981 4.14185 4.22461
74Ge GCN2850 3.98664 4.14437 4.23118
JUN45 3.98951 4.16453 4.24903
jj44b 3.98955 4.17113 4.25434
76Ge GCN2850 4.00136 4.20041 4.28574
JUN45 4.00456 4.21257 4.29678
jj44b 4.00422 4.21636 4.29966
127I B&D 4.6651 4.89861 4.97959
SN100PN 4.66554 4.92354 4.99969
Trun. SN100PN 4.66559 4.92575 5.00184
GCN5082 4.66544 4.91281 4.98982
GCN5082* 4.66425 4.91304 4.98979
133Cs SN100PN 4.72559 4.99394 5.06814
SN100PN* 4.7252 4.99335 5.06752
GCN5082 4.72545 4.99188 5.06646
GCN5082* 4.72493 4.99011 5.06455
Appendix B Comparison to nuclear observables
Here we compare experimental nuclear observables to the theoretical shell model values, namely the nuclear energy levels, as well as the magnetic dipole and electric quadrupole moments and transitions. For each of these observables, we provide the values for 40Ar, 70,72,74,76Ge and 133Cs below, whilst the 73Ge and 127I comparisons can be found in [39].
B.1 Argon-40
Figure 5: Experimental energy levels (keV) for 40Ar compared with levels from the SDPF-NR, SDPF-U, SDPF-MU and EPQQM shell model calculations.
Table 8: Magnetic dipole and electric quadrupole moments, and electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions, between low-lying states of 40Ar. Experimental transition values from [56], while moments taken from [57, 58].
Q [$e$fm2] $\mu$ [n.m.]
State SDPF-NR SDPF-U SDPF-MU EPQQM Exp. SDPF-NR SDPF-U SDPF-MU EPQQM Exp.
$2_{1}^{+}$ 8.30 9.01 10.59 -0.27 +1(4) -0.401 -0.456 -0.741 0.103 -0.04(6)
B(E2) [$e^{2}$fm4] B(M1) [n.m.2]
$2_{1}^{+}\rightarrow 0_{\text{gs}}^{+}$ 50.31 48.86 46.33 59.58 73.13(325)
$2_{2}^{+}\rightarrow 0_{\text{gs}}^{+}$ 8.024 10.40 7.671 7.294 9.670(1463)
$0_{2}^{+}\rightarrow 2_{1}^{+}$ 6.980 14.50 14.68 0.03158 43.07(650)
$2_{2}^{+}\rightarrow 2_{1}^{+}$ 23.76 17.83 28.13 60.55 146.26(4063) 0.2710 0.3055 0.2952 0.1203 0.06623(1074)
B.2 Germanium-70,72,74,76
Figure 6: Low-lying experimental energy levels (keV) for germanium isotopes compared with levels from shell model calculations using the JUN45 and jj44b interactions.
Table 9: Magnetic dipole and electric quadrupole moments, for low-lying states of 70,72,74,76Ge. Experimental moments taken from [57, 58, 59].
70Ge 72Ge
Q [$e$fm2] $\mu$ [n.m.] Q [$e$fm2] $\mu$ [n.m.]
State JUN45 jj44b Exp. JUN45 jj44b Exp. JUN45 jj44b Exp. JUN45 jj44b Exp.
$2_{1}^{+}$ +14.32 +21.35 +3(6) +0.72 +0.59 +0.64(10) +18.30 +15.70 -13(6) +0.62 +0.52 +0.56(8)
$2_{2}^{+}$18.63 -21.70 +1.3 +1.1 +1.0(4) -19.09 -16.14 +1.4 +1.0 +0.5(3)
$4_{1}^{+}$ +4.03 +6.48 +1.32 +1.12 +0.88(124) +13.53 +5.98 +1.12 +0.67 +1.0(3)
74Ge 76Ge
$2_{1}^{+}$ +16.99 -7.22 -19(2) +0.59 +0.61 +0.56(8) +2.99 -18.81 -19(6) +0.737 +0.596 +0.53(8)
$2_{2}^{+}$16.13 +6.67 -26(6) +1.25 +0.96 +0.75(15) -0.72 +20.20 +1.16 +1.07 +0.64(10)
$4_{1}^{+}$ +16.89 -10.15 +0.81 +0.91 +1.3(4) -0.76 -18.06 +1.20 +0.87 +0.8(6)
Table 10: Electric quadrupole [B(E2)] transitions between low-lying states of 70,72,74,76Ge. Experimental transitions taken from [60, 61, 62, 63].
70Ge 72Ge 74Ge 76Ge
B(E2) [$e^{2}$fm4] B(E2) [$e^{2}$fm4] B(E2) [$e^{2}$fm4] B(E2) [$e^{2}$fm4]
Transition JUN45 jj44b Exp. JUN45 jj44b Exp. JUN45 jj44b Exp. JUN45 jj44b Exp.
$2_{1}^{+}\rightarrow 0_{\text{gs}}^{+}$ 437.3 602.4 356.4(69) 462.6 632.0 418.1(71) 542.8 648.3 609.0(74) 544.8 608.0 550.9(40)
$0_{2}^{+}\rightarrow 2_{1}^{+}$ 475.9 104.7 822.5(1199)
[…]
$2_{1}^{+}\rightarrow 0_{2}^{+}$ 116.0 25.3 316.7(53)
$4_{1}^{+}\rightarrow 2_{1}^{+}$ 772.7 885.4 756.6(554) 738.3 809.7 698.0(153)
B.3 Cesium-133
Figure 7: Experimental energy levels (keV) for 133Cs, compared with levels from SN100PN and GCN5082 shell model calculations for two different valence space truncations.
Table 11: Magnetic dipole and electric quadrupole moments for low-lying states of 133Cs. Experimental moments taken from [64, 57, 58]. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
Q [$e$fm2] $\mu$ [n.m.]
State SN100PN SN100PN* GCN5082 GCN5082* Exp. SN100PN SN100PN* GCN5082 GCN5082* Exp.
$7/2_{\text{gs}}^{+}$14.11 -13.28 -18.71 -17.13 -0.343(10) 2.6649 2.6880 2.6715 2.7222 +2.5778(14)
$5/2_{1}^{+}$66.40 -67.47 -63.36 -60.36 30(2) [^1] Sign not specified 3.54 3.62 3.36 3.50 +3.45(2)
$5/2_{2}^{+}$ 2.22 2.19 2.38 2.33 +2.25(15)
Table 12: Electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions between low-lying states of 133Cs. Experimental transition values from [65]. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
B(E2) [$e^{2}$fm4]
Transition SN100PN SN100PN* GCN5082 GCN5082* Exp.
$5/2_{1}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 178.7 155.2 242.8 177.2 233.9(161)
$5/2_{2}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 1772 1863 1715 1752 1153.3(726)
$3/2_{1}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 1005 733.5 525.5 217.6 483.9(1210)
B(M1) [n.m.2]
$5/2_{1}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 0.000006111 0.00008033 0.0003863 0.0001393 0.00426(4)
$5/2_{2}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 0.01032 0.004174 0.01096 0.004590 0.00226(14)
Appendix C EFT/Nuclear Operators
The nuclear operators employed in Eq. (IV.1) have the form
$\begin{split}M_{JM}(q\vec{x})&\equiv j_{J}(qx)Y_{JM}(\Omega_{x}),\\ \Sigma^{\prime}_{JM}(q\vec{x})&\equiv-i\left[\frac{\vec{\nabla}}{q}\times\vec{M}^{M}_{JJ}(q\vec{x})\right]\cdot\vec{\sigma}_{N}=[J]^{-1}\left[-\sqrt{J}\ \vec{M}^{M}_{JJ+1}(q\vec{x})+\sqrt{J+1}\ \vec{M}^{M}_{JJ-1}(q\vec{x})\right]\cdot\vec{\sigma}_{N},\\ \Phi^{\prime\prime}_{JM}(q\vec{x})&\equiv i\left(\frac{\vec{\nabla}}{q}M_{JM}(q\vec{x})\right)\cdot\left(\vec{\sigma}_{N}\times\frac{1}{q}\vec{\nabla}\right),\end{split}$ (33)
where   $\vec{M}_{JL}^{M}\equiv j_{L}(qx)\vec{Y}_{JLM}$, $[J]=\sqrt{2J+1}$, and $\vec{\sigma}_{N}$ is the nucleon spin operator. Here, $j_{J}(qx)$ is the spherical Bessel function, $Y_{JM}$ is the spherical harmonic, and $\vec{Y}_{JLM}$ is the vector spherical harmonic.
The Wigner-Eckart theorem has been used to express matrix elements in reduced form, using the convention
$\begin{split}&\langle j^{\prime}m^{\prime}|T_{JM}|jm\rangle=(-1)^{j^{\prime}-m^{\prime}}\begin{pmatrix}\vspace{2mm}j^{\prime}&J&j\\ \vspace{2mm}-m^{\prime}&M&m\hskip 5.69054pt\end{pmatrix}\langle j^{\prime}||T_{J}||j\rangle.\end{split}$ (34)
The one-nucleon matrix elements in Eq. (IV.1) have analytic expressions of the form [15]:
$\Big{\langle}n^{\prime}l^{\prime}\frac{1}{2}j^{\prime}\Big{|}\Big{|}M_{J}\Big{|}\Big{|}nl\frac{1}{2}j\Big{\rangle}$ $=\langle n^{\prime}l^{\prime}|j_{J}(qr_{i})|nl\rangle(-1)^{j+1/2+J}\sqrt{\frac{1}{4\pi}}\,\bigl{[}(2j^{\prime}+1)(2j+1)\bigr{]}^{\frac{1}{2}}\bigl{[}(2J+1)(2l+1)(2l^{\prime}+1)\bigr{]}^{\frac{1}{2}}$
×$\quad\times\left(\begin{array}[]{ccc}l^{\prime}&J&l\\ 0&0&0\end{array}\right)\left\{\begin{matrix}l^{\prime}&j^{\prime}&\nicefrac{{1}}{{2}}\\ j&l&J\end{matrix}\right\},$ (37)
$\Big{\langle}n^{\prime}l^{\prime}\frac{1}{2}j^{\prime}\Big{|}\Big{|}j_{J^{\prime}}(pr_{i})\bigl{[}Y_{J^{\prime}}(\hat{\mathbf{r}}_{i})\bm{\sigma}_{i}\bigr{]}^{J}\Big{|}\Big{|}nl\frac{1}{2}j\Big{\rangle}$ $=\langle n^{\prime}l^{\prime}|j_{J^{\prime}}(qr_{i})|nl\rangle(-1)^{l^{\prime}}\sqrt{\frac{6}{4\pi}}\,\bigl{[}(2l^{\prime}+1)(2l+1)(2j^{\prime}+1)(2j+1)\bigr{]}^{\frac{1}{2}}$
×$\times\bigl{[}(2J^{\prime}+1)(2J+1)\bigr{]}^{\frac{1}{2}}\left(\begin{matrix}l^{\prime}&J^{\prime}&l\\ 0&0&0\end{matrix}\right)\left\{\begin{matrix}l^{\prime}&l&J^{\prime}\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&1\\ j^{\prime}&j&J\end{matrix}\right\},$ (38)
$\langle n^{\prime}l^{\prime}j^{\prime}||\Phi^{\prime\prime}_{J}||nlj\rangle$ $=(-1)^{l^{\prime}}\frac{6}{\sqrt{4\pi}}\sqrt{(2j^{\prime}+1)(2j+1)(2l^{\prime}+1)}$ (39)
×$\times\left\{\sqrt{(J+1)(2J+3)}\sum_{L=J}^{J+1}(-1)^{J+L}(2L+1)\Bigg{\{}\begin{matrix}J+1&1&L\\ 1&J&1\end{matrix}\right\}\left\{\begin{matrix}l^{\prime}&l&L\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&1\\ j^{\prime}&j&J\end{matrix}\right\}$
×$\quad\times\bigg{[}\sqrt{(l+1)(2l+3)}\left\{\begin{matrix}J+1&1&L\\ l&l^{\prime}&l+1\end{matrix}\right\}\left(\begin{matrix}l^{\prime}&J+1&l+1\\ 0&0&0\end{matrix}\right)\Big{\langle}n^{\prime}l^{\prime}\Big{|}j_{J+1}(qr_{i})\left(\frac{\partial}{\partial(qr_{i})}-\frac{l}{qr_{i}}\right)\Big{|}nl\Big{\rangle}$
$\quad\quad-\sqrt{l(2l-1)}\left\{\begin{matrix}J+1&1&L\\ l&l^{\prime}&l-1\end{matrix}\right\}\left(\begin{matrix}l^{\prime}&J+1&l-1\\ 0&0&0\end{matrix}\right)\Big{\langle}n^{\prime}l^{\prime}\Big{|}j_{J+1}(qr_{i})\left(\frac{\partial}{\partial(qr_{i})}+\frac{l+1}{qr_{i}}\right)\Big{|}nl\Big{\rangle}\bigg{]}$
$+\sqrt{J(2J-1)}\sum_{L=J-1}^{J}(-1)^{J+L}(2L+1)\left\{\begin{matrix}J-1&1&L\\ 1&J&1\end{matrix}\right\}\left\{\begin{matrix}l^{\prime}&l&L\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&1\\ j^{\prime}&j&J\end{matrix}\right\}$
×$\quad\times\bigg{[}\sqrt{(l+1)(2l+3)}\left\{\begin{matrix}J-1&1&L\\ l&l^{\prime}&l+1\end{matrix}\right\}\left(\begin{matrix}l^{\prime}&J-1&l+1\\ 0&0&0\end{matrix}\right)\Big{\langle}n^{\prime}l^{\prime}\Big{|}j_{J-1}(qr_{i})\left(\frac{\partial}{\partial(qr_{i})}-\frac{l}{qr_{i}}\right)\Big{|}nl\Big{\rangle}$
$\qquad\quad-\sqrt{l(2l-1)}\left\{\begin{matrix}J-1&1&L\\ l&l^{\prime}&l-1\end{matrix}\right\}\left(\begin{matrix}l^{\prime}&J-1&l-1\\ 0&0&0\end{matrix}\right)\Big{\langle}n^{\prime}l^{\prime}\Big{|}j_{J-1}(qr_{i})\left(\frac{\partial}{\partial(qr_{i})}+\frac{l+1}{qr_{i}}\right)\Big{|}nl\Big{\rangle}\bigg{]}\Bigg{\}}.$
Appendix D Form factors
In this section we plot the form factors for the individual responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ using the shell model calculations performed in this work, where the $\Sigma^{\prime}$ response is only plotted for isotopes with non-zero nuclear ground state spin. This is presented for 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs, highlighting the corresponding COHERENT range of interest for each of these nuclei. The form factor expressions employed to plot these are presented in [39], except for 133Cs which is provided in Appendix D. The KN and SHF form factors were plotted using $(ZF\left(q,r_{\text{rms}}^{p})+NF(q,r_{\text{rms}}^{n})\right)^{2}$, while the shell model counterparts were plotted using $\sum_{N,N^{\prime}=p,n}F^{(N,N^{\prime})}_{X,X}(q^{2})$, where $F^{(N,N^{\prime})}_{X,X}(q^{2})$ is defined in Eq. (40).
D.1 Argon-40
Figure 8: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}={r^{p}_{\text{rms}}}^{+0.2}_{-0}$. The COHERENT region-of-interest is also displayed.
Figure 9: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.2 Germanium-70
Figure 10: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 11: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.3 Germanium-72
Figure 12: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 13: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.4 Germanium-73
Figure 14: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 15: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
Figure 16: $\Sigma^{\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.5 Germanium-74
Figure 17: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 18: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.6 Germanium-76
Figure 19: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 20: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.7 Iodine-127
Figure 21: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 22: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
Figure 23: $\Sigma^{\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
D.8 Cesium-133
Figure 24: SI ($M$) form factors employing the different shell model interactions used for this isotope. This response is compared against the KN and SHF distributions, where the uncertainty bands stem from the $r^{n}_{\text{rms}}$ uncertainty. The COHERENT range-of-interest is also displayed.
Figure 25: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
Figure 26: $\Sigma^{\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
133Cs Form Factor Expressions
The explicit nuclear form factors for 133Cs are given below using the convention of [35]:
$F^{(N,N^{\prime})}_{X,Y}(q^{2})\equiv\frac{4\pi}{2J_{i}+1}\sum_{J=0}^{2J_{i}}\langle J_{i}||X_{J}^{(N)}||J_{i}\rangle\langle J_{i}||Y^{(N^{\prime})}_{J}||J_{i}\rangle,$ (40)
and can be used to obtain the $\mathcal{F}_{N}^{Y}(q^{2})$ in Eq. (IV.1). Since no interference occurs between operators in CEvNS, we simplify the notation and write $F^{(N,N^{\prime})}_{X,X}(q^{2})\equiv F^{(N,N^{\prime})}_{X}(q^{2})$.
SN100PN
$F_{M}^{(p,p)}=e^{-2y}(3020.-11000.y+15500.y^{2}-10800.y^{3}+4130.y^{4}-881.y^{5}+103.y^{6}-6.y^{7}+0.145y^{8}-0.000591y^{9}+6.62\times 10^{-7}y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,p)}=e^{-2y}(0.204-0.18y+0.0841y^{2}+0.168y^{3}-0.0456y^{4}-0.00246y^{5}+0.014y^{6}-0.00278y^{7}+0.000806y^{8}+1.9\times 10^{-8}y^{9}+1.19\times 10^{-12}y^{10})$
$F_{\Sigma^{\prime}}^{(p,p)}=e^{-2y}(0.408-3.09y+8.18y^{2}-9.16y^{3}+5.23y^{4}-1.62y^{5}+0.287y^{6}-0.0273y^{7}+0.00167y^{8}+1.8\times 10^{-7}y^{9}+2.77\times 10^{-11}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(76.-185.y+171.y^{2}-75.9y^{3}+17.4y^{4}-2.01y^{5}+0.101y^{6}-0.0012y^{7}+4.14\times 10^{-6}y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.00327+0.00151y-0.00238y^{2}-0.000124y^{3}+0.00104y^{4}-0.00025y^{5}+0.000026y^{6}-4.22\times 10^{-9}y^{7}+2.11\times 10^{-9}y^{8})$
$F_{\Delta}^{(p,p)}=e^{-2y}(2.73-6.56y+6.44y^{2}-2.97y^{3}+0.704y^{4}-0.0813y^{5}+0.00378y^{6}-1.37\times 10^{-6}y^{7}+1.91\times 10^{-10}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(-479.+1460.y-1680.y^{2}+948.y^{3}-284.y^{4}+45.8y^{5}-3.72y^{6}+0.127y^{7}-0.000979y^{8}+1.66\times 10^{-6}y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,p)}=e^{-2y}(1.06-5.26y+8.26y^{2}-6.02y^{3}+2.24y^{4}-0.451y^{5}+0.0461y^{6}-0.00218y^{7}+1.81\times 10^{-8}y^{8}+7.25\times 10^{-11}y^{9})$
$F_{M}^{(p,n)}=e^{-2y}(4280.-17000.y+26500.y^{2}-20900.y^{3}+9210.y^{4}-2350.y^{5}+342.y^{6}-27.y^{7}+1.y^{8}-0.0128y^{9}+0.000025y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,n)}=e^{-2y}(-0.0048+0.0153y-0.0323y^{2}+0.0186y^{3}+0.000919y^{4}-0.00533y^{5}+0.00219y^{6}-0.000397y^{7}+0.0000337y^{8}-9.97\times 10^{-7}y^{9}+1.92\times 10^{-11}y^{10})$
$F_{\Sigma^{\prime}}^{(p,n)}=e^{-2y}(-0.00961+0.0804y-0.264y^{2}+0.382y^{3}-0.279y^{4}+0.112y^{5}-0.0259y^{6}+0.0034y^{7}-0.000232y^{8}+6.38\times 10^{-6}y^{9}+8.41\times 10^{-10}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(161.-449.y+478.y^{2}-252.y^{3}+70.9y^{4}-10.8y^{5}+0.831y^{6}-0.0265y^{7}+0.000156y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,n)}=e^{-2y}(-0.0171+0.0166y+0.00245y^{2}-0.0105y^{3}+0.00504y^{4}-0.00096y^{5}+0.000072y^{6}-1.28\times 10^{-6}y^{7}-2.6\times 10^{-8}y^{8})$
$F_{\Delta}^{(p,n)}=e^{-2y}(0.199-0.526y+0.584y^{2}-0.324y^{3}+0.0954y^{4}-0.0148y^{5}+0.0011y^{6}-0.0000299y^{7}+6.02\times 10^{-9}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(-1020.+3440.y-4510.y^{2}+2950.y^{3}-1060.y^{4}+213.y^{5}-23.6y^{6}+1.31y^{7}-0.0295y^{8}+0.0000625y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,n)}=e^{-2y}(0.0768-0.401y+0.694y^{2}-0.581y^{3}+0.26y^{4}-0.0643y^{5}+0.00865y^{6}-0.000584y^{7}+0.0000155y^{8}+2.25\times 10^{-9}y^{9})$
$F_{M\Phi^{\prime\prime}}^{(n,p)}=e^{-2y}(-679.+2290.y-2960.y^{2}+1920.y^{3}-680.y^{4}+135.y^{5}-14.5y^{6}+0.749y^{7}-0.0132y^{8}+0.0000625y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.0249+0.144y-0.304y^{2}+0.29y^{3}-0.141y^{4}+0.0379y^{5}-0.00559y^{6}+0.000415y^{7}-0.0000118y^{8}+2.25\times 10^{-9}y^{9})$
$F_{M}^{(n,n)}=e^{-2y}(6070.-26100.y+44600.y^{2}-39200.y^{3}+19700.y^{4}-5870.y^{5}+1040.y^{6}-106.y^{7}+5.58y^{8}-0.124y^{9}+0.000948y^{10})$
$F_{\Sigma^{\prime\prime}}^{(n,n)}=e^{-2y}(0.000113-0.00062y+0.00352y^{2}-0.00581y^{3}+0.00463y^{4}-0.00205y^{5}+0.00053y^{6}-0.0000795y^{7}+6.33\times 10^{-6}y^{8}-2.1\times 10^{-7}y^{9}+2.77\times 10^{-9}y^{10})$
$F_{\Sigma^{\prime}}^{(n,n)}=e^{-2y}(0.000226-0.00207y+0.00999y^{2}-0.018y^{3}+0.0157y^{4}-0.00743y^{5}+0.00209y^{6}-0.000355y^{7}+0.0000357y^{8}-1.84\times 10^{-6}y^{9}+3.92\times 10^{-8}y^{10})$
$F_{\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(343.-1070.y+1300.y^{2}-797.y^{3}+269.y^{4}-51.4y^{5}+5.41y^{6}-0.287y^{7}+0.00593y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(n,n)}=e^{-2y}(0.0891-0.215y+0.223y^{2}-0.122y^{3}+0.038y^{4}-0.00687y^{5}+0.000686y^{6}-0.0000332y^{7}+6.04\times 10^{-7}y^{8})$
$F_{\Delta}^{(n,n)}=e^{-2y}(0.0144-0.0417y+0.0517y^{2}-0.0334y^{3}+0.012y^{4}-0.00238y^{5}+0.000252y^{6}-0.0000128y^{7}+2.54\times 10^{-7}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(-1440.+5360.y-7780.y^{2}+5760.y^{3}-2400.y^{4}+578.y^{5}-80.2y^{6}+6.04y^{7}-0.213y^{8}+0.00237y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,n)}=e^{-2y}(-0.00181+0.0109y-0.0246y^{2}+0.0263y^{3}-0.015y^{4}+0.00482y^{5}-0.000901y^{6}+0.0000936y^{7}-4.82\times 10^{-6}y^{8}+9.94\times 10^{-8}y^{9})$
SN100PN Expanded
$F_{M}^{(p,p)}=e^{-2y}(3020.-11000.y+15500.y^{2}-10800.y^{3}+4120.y^{4}-876.y^{5}+102.y^{6}-5.91y^{7}+0.143y^{8}-0.000608y^{9}+7.17\times 10^{-7}y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,p)}=e^{-2y}(0.206-0.19y+0.1y^{2}+0.166y^{3}-0.0494y^{4}+0.000185y^{5}+0.0135y^{6}-0.00271y^{7}+0.000818y^{8}+2.12\times 10^{-8}y^{9}+1.62\times 10^{-12}y^{10})$
$F_{\Sigma^{\prime}}^{(p,p)}=e^{-2y}(0.412-3.11y+8.21y^{2}-9.17y^{3}+5.21y^{4}-1.61y^{5}+0.283y^{6}-0.0269y^{7}+0.00167y^{8}+2.15\times 10^{-7}y^{9}+3.99\times 10^{-11}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(75.3-184.y+170.y^{2}-75.5y^{3}+17.3y^{4}-2.01y^{5}+0.102y^{6}-0.00125y^{7}+4.48\times 10^{-6}y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.0028+0.00224y-0.00221y^{2}-0.000663y^{3}+0.00134y^{4}-0.000307y^{5}+0.0000285y^{6}+1.9\times 10^{-8}y^{7}+2.31\times 10^{-9}y^{8})$
$F_{\Delta}^{(p,p)}=e^{-2y}(2.77-6.65y+6.53y^{2}-3.01y^{3}+0.713y^{4}-0.0822y^{5}+0.00382y^{6}-1.66\times 10^{-6}y^{7}+2.77\times 10^{-10}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(-477.+1450.y-1670.y^{2}+944.y^{3}-283.y^{4}+45.6y^{5}-3.71y^{6}+0.126y^{7}-0.00101y^{8}+1.79\times 10^{-6}y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,p)}=e^{-2y}(1.07-5.31y+8.33y^{2}-6.06y^{3}+2.25y^{4}-0.452y^{5}+0.046y^{6}-0.00217y^{7}+1.65\times 10^{-8}y^{8}+1.05\times 10^{-10}y^{9})$
$F_{M}^{(p,n)}=e^{-2y}(4280.-17000.y+26500.y^{2}-20800.y^{3}+9180.y^{4}-2330.y^{5}+339.y^{6}-26.7y^{7}+0.998y^{8}-0.013y^{9}+0.0000266y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,n)}=e^{-2y}(-0.0042+0.0114y-0.0194y^{2}+0.00958y^{3}+0.00118y^{4}-0.00307y^{5}+0.00117y^{6}-0.000213y^{7}+0.00002y^{8}-7.23\times 10^{-7}y^{9}+1.64\times 10^{-11}y^{10})$
$F_{\Sigma^{\prime}}^{(p,n)}=e^{-2y}(-0.00841+0.0695y-0.22y^{2}+0.309y^{3}-0.219y^{4}+0.0848y^{5}-0.0189y^{6}+0.00241y^{7}-0.000163y^{8}+4.55\times 10^{-6}y^{9}+7.28\times 10^{-10}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(161.-449.y+479.y^{2}-252.y^{3}+70.9y^{4}-10.8y^{5}+0.84y^{6}-0.0273y^{7}+0.000166y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,n)}=e^{-2y}(-0.0115+0.00924y+0.00526y^{2}-0.00957y^{3}+0.00411y^{4}-0.00075y^{5}+0.0000598y^{6}-1.43\times 10^{-6}y^{7}-2.15\times 10^{-8}y^{8})$
$F_{\Delta}^{(p,n)}=e^{-2y}(0.18-0.468y+0.513y^{2}-0.285y^{3}+0.0835y^{4}-0.0127y^{5}+0.000896y^{6}-0.0000218y^{7}+5.22\times 10^{-9}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(-1020.+3460.y-4530.y^{2}+2960.y^{3}-1060.y^{4}+213.y^{5}-23.6y^{6}+1.32y^{7}-0.0299y^{8}+0.0000666y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,n)}=e^{-2y}(0.0695-0.359y+0.608y^{2}-0.506y^{3}+0.226y^{4}-0.0552y^{5}+0.00718y^{6}-0.000457y^{7}+0.000011y^{8}+1.95\times 10^{-9}y^{9})$
$F_{M\Phi^{\prime\prime}}^{(n,p)}=e^{-2y}(-675.+2280.y-2950.y^{2}+1910.y^{3}-677.y^{4}+134.y^{5}-14.5y^{6}+0.752y^{7}-0.0136y^{8}+0.0000666y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.0218+0.124y-0.253y^{2}+0.234y^{3}-0.11y^{4}+0.0286y^{5}-0.00409y^{6}+0.000299y^{7}-8.57\times 10^{-6}y^{8}+1.95\times 10^{-9}y^{9})$
$F_{M}^{(n,n)}=e^{-2y}(6060.-26100.y+44600.y^{2}-39200.y^{3}+19600.y^{4}-5840.y^{5}+1030.y^{6}-105.y^{7}+5.58y^{8}-0.127y^{9}+0.000994y^{10})$
$F_{\Sigma^{\prime\prime}}^{(n,n)}=e^{-2y}(0.0000858-0.000388y+0.00137y^{2}-0.00192y^{3}+0.00139y^{4}-0.000568y^{5}+0.000138y^{6}-0.0000202y^{7}+1.72\times 10^{-6}y^{8}-7.53\times 10^{-8}y^{9}+1.43\times 10^{-9}y^{10})$
$F_{\Sigma^{\prime}}^{(n,n)}=e^{-2y}(0.000172-0.00154y+0.0061y^{2}-0.0102y^{3}+0.0086y^{4}-0.00401y^{5}+0.00112y^{6}-0.000189y^{7}+0.0000185y^{8}-9.03\times 10^{-7}y^{9}+2.03\times 10^{-8}y^{10})$
$F_{\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(344.-1080.y+1310.y^{2}-803.y^{3}+270.y^{4}-51.5y^{5}+5.45y^{6}-0.293y^{7}+0.00621y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(n,n)}=e^{-2y}(0.0471-0.113y+0.115y^{2}-0.0603y^{3}+0.0179y^{4}-0.00312y^{5}+0.000316y^{6}-0.0000171y^{7}+3.82\times 10^{-7}y^{8})$
$F_{\Delta}^{(n,n)}=e^{-2y}(0.0117-0.0328y+0.0397y^{2}-0.0256y^{3}+0.00922y^{4}-0.0018y^{5}+0.000179y^{6}-7.73\times 10^{-6}y^{7}+1.32\times 10^{-7}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(-1440.+5380.y-7820.y^{2}+5780.y^{3}-2400.y^{4}+577.y^{5}-80.2y^{6}+6.07y^{7}-0.217y^{8}+0.00248y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,n)}=e^{-2y}(-0.00142+0.00836y-0.0178y^{2}+0.0182y^{3}-0.0101y^{4}+0.00319y^{5}-0.000576y^{6}+0.0000566y^{7}-2.65\times 10^{-6}y^{8}+5.16\times 10^{-8}y^{9})$
GCN5082
$F_{M}^{(p,p)}=e^{-2y}(3020.-11000.y+15500.y^{2}-10900.y^{3}+4220.y^{4}-914.y^{5}+109.y^{6}-6.58y^{7}+0.165y^{8}-0.000663y^{9}+7.35\times 10^{-7}y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,p)}=e^{-2y}(0.185-0.0951y-0.0302y^{2}+0.167y^{3}+0.0107y^{4}-0.0318y^{5}+0.0186y^{6}-0.00298y^{7}+0.000747y^{8}+1.84\times 10^{-8}y^{9}+2.34\times 10^{-12}y^{10})$
$F_{\Sigma^{\prime}}^{(p,p)}=e^{-2y}(0.37-2.86y+7.74y^{2}-8.75y^{3}+5.02y^{4}-1.57y^{5}+0.281y^{6}-0.0275y^{7}+0.00167y^{8}+3.2\times 10^{-7}y^{9}+7.21\times 10^{-11}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(92.1-225.y+208.y^{2}-93.y^{3}+21.5y^{4}-2.52y^{5}+0.128y^{6}-0.00143y^{7}+4.59\times 10^{-6}y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.0147-0.00888y+0.000482y^{2}+0.00251y^{3}-(3.38*10^{(}-6))y^{4}-0.0000492y^{5}+0.0000216y^{6}+4.51\times 10^{-7}y^{7}+3.47\times 10^{-9}y^{8})$
$F_{\Delta}^{(p,p)}=e^{-2y}(2.65-6.37y+6.21y^{2}-2.86y^{3}+0.678y^{4}-0.0785y^{5}+0.00367y^{6}-2.32\times 10^{-6}y^{7}+5.1\times 10^{-10}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(-528.+1600.y-1850.y^{2}+1050.y^{3}-320.y^{4}+52.3y^{5}-4.34y^{6}+0.15y^{7}-0.00112y^{8}+1.84\times 10^{-6}y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,p)}=e^{-2y}(0.991-5.02y+7.93y^{2}-5.78y^{3}+2.16y^{4}-0.436y^{5}+0.045y^{6}-0.00214y^{7}+3.71\times 10^{-8}y^{8}+1.91\times 10^{-10}y^{9})$
$F_{M}^{(p,n)}=e^{-2y}(4280.-17000.y+26500.y^{2}-21000.y^{3}+9290.y^{4}-2380.y^{5}+351.y^{6}-28.1y^{7}+1.06y^{8}-0.0135y^{9}+0.0000258y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,n)}=e^{-2y}(-0.00421+0.0148y-0.029y^{2}+0.0133y^{3}+0.00321y^{4}-0.00556y^{5}+0.0021y^{6}-0.000387y^{7}+0.0000362y^{8}-1.17\times 10^{-6}y^{9}+2.84\times 10^{-11}y^{10})$
$F_{\Sigma^{\prime}}^{(p,n)}=e^{-2y}(-0.00841+0.0722y-0.243y^{2}+0.357y^{3}-0.265y^{4}+0.108y^{5}-0.0256y^{6}+0.00346y^{7}-0.000246y^{8}+7.12\times 10^{-6}y^{9}+1.42\times 10^{-9}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(175.-486.y+519.y^{2}-274.y^{3}+77.4y^{4}-11.8y^{5}+0.916y^{6}-0.0292y^{7}+0.000162y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,n)}=e^{-2y}(-0.0389+0.0585y-0.0328y^{2}+0.00374y^{3}+0.00233y^{4}-0.000795y^{5}+0.0000795y^{6}-1.68\times 10^{-6}y^{7}-4.12\times 10^{-8}y^{8})$
$F_{\Delta}^{(p,n)}=e^{-2y}(0.223-0.588y+0.643y^{2}-0.351y^{3}+0.102y^{4}-0.0157y^{5}+0.00117y^{6}-0.0000316y^{7}+1.03\times 10^{-8}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(-1000.+3390.y-4440.y^{2}+2920.y^{3}-1050.y^{4}+214.y^{5}-23.9y^{6}+1.35y^{7}-0.031y^{8}+0.0000646y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,n)}=e^{-2y}(0.0832-0.441y+0.762y^{2}-0.632y^{3}+0.28y^{4}-0.0688y^{5}+0.00921y^{6}-0.000627y^{7}+0.000017y^{8}+3.83\times 10^{-9}y^{9})$
$F_{M\Phi^{\prime\prime}}^{(n,p)}=e^{-2y}(-747.+2520.y-3260.y^{2}+2120.y^{3}-753.y^{4}+150.y^{5}-16.3y^{6}+0.839y^{7}-0.0144y^{8}+0.0000646y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.0225+0.133y-0.283y^{2}+0.272y^{3}-0.134y^{4}+0.0366y^{5}-0.00553y^{6}+0.000422y^{7}-0.0000125y^{8}+3.83\times 10^{-9}y^{9})$
$F_{M}^{(n,n)}=e^{-2y}(6070.-26100.y+44500.y^{2}-39200.y^{3}+19700.y^{4}-5880.y^{5}+1040.y^{6}-106.y^{7}+5.58y^{8}-0.122y^{9}+0.000919y^{10})$
$F_{\Sigma^{\prime\prime}}^{(n,n)}=e^{-2y}(0.0000957-0.000626y+0.0036y^{2}-0.00574y^{3}+0.00447y^{4}-0.00195y^{5}+0.000504y^{6}-0.0000768y^{7}+6.4\times 10^{-6}y^{8}-2.28\times 10^{-7}y^{9}+3.28\times 10^{-9}y^{10})$
$F_{\Sigma^{\prime}}^{(n,n)}=e^{-2y}(0.000191-0.0018y+0.00906y^{2}-0.0167y^{3}+0.0148y^{4}-0.00718y^{5}+0.00205y^{6}-0.000357y^{7}+0.0000371y^{8}-1.98\times 10^{-6}y^{9}+4.34\times 10^{-8}y^{10})$
$F_{\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(333.-1040.y+1260.y^{2}-773.y^{3}+261.y^{4}-49.7y^{5}+5.23y^{6}-0.277y^{7}+0.00574y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(n,n)}=e^{-2y}(0.103-0.247y+0.254y^{2}-0.137y^{3}+0.0425y^{4}-0.00762y^{5}+0.000756y^{6}-0.0000363y^{7}+6.56\times 10^{-7}y^{8})$
$F_{\Delta}^{(n,n)}=e^{-2y}(0.0187-0.0538y+0.0652y^{2}-0.0412y^{3}+0.0145y^{4}-0.00283y^{5}+0.000293y^{6}-0.0000145y^{7}+2.79\times 10^{-7}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(-1420.+5270.y-7660.y^{2}+5670.y^{3}-2360.y^{4}+569.y^{5}-79.y^{6}+5.94y^{7}-0.209y^{8}+0.0023y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,n)}=e^{-2y}(-0.00189+0.0116y-0.026y^{2}+0.0276y^{3}-0.0157y^{4}+0.00509y^{5}-0.000965y^{6}+0.000102y^{7}-5.31\times 10^{-6}y^{8}+1.1\times 10^{-7}y^{9})$
GCN5082 Expanded
$F_{M}^{(p,p)}=e^{-2y}(3020.-11000.y+15500.y^{2}-10900.y^{3}+4200.y^{4}-907.y^{5}+108.y^{6}-6.46y^{7}+0.162y^{8}-0.000701y^{9}+8.42\times 10^{-7}y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,p)}=e^{-2y}(0.193-0.121y+0.00697y^{2}+0.175y^{3}-0.00996y^{4}-0.0217y^{5}+0.0172y^{6}-0.0029y^{7}+0.000786y^{8}+3.71\times 10^{-8}y^{9}+3.42\times 10^{-12}y^{10})$
$F_{\Sigma^{\prime}}^{(p,p)}=e^{-2y}(0.386-2.97y+7.97y^{2}-8.97y^{3}+5.13y^{4}-1.6y^{5}+0.284y^{6}-0.0276y^{7}+0.0017y^{8}+3.15\times 10^{-7}y^{9}+8.5\times 10^{-11}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(91.3-223.y+207.y^{2}-92.9y^{3}+21.6y^{4}-2.55y^{5}+0.131y^{6}-0.00154y^{7}+5.26\times 10^{-6}y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.015-0.00813y+0.000808y^{2}+0.00263y^{3}-(8.21*10^{(}-6))y^{4}-0.0000477y^{5}+0.0000242y^{6}+5.25\times 10^{-7}y^{7}+4.03\times 10^{-9}y^{8})$
$F_{\Delta}^{(p,p)}=e^{-2y}(2.74-6.58y+6.44y^{2}-2.97y^{3}+0.703y^{4}-0.0812y^{5}+0.00379y^{6}-2.48\times 10^{-6}y^{7}+5.93\times 10^{-10}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(-525.+1600.y-1850.y^{2}+1050.y^{3}-319.y^{4}+52.3y^{5}-4.34y^{6}+0.151y^{7}-0.00118y^{8}+2.1\times 10^{-6}y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,p)}=e^{-2y}(1.03-5.18y+8.17y^{2}-5.96y^{3}+2.22y^{4}-0.447y^{5}+0.046y^{6}-0.00218y^{7}+2.02\times 10^{-8}y^{8}+2.24\times 10^{-10}y^{9})$
$F_{M}^{(p,n)}=e^{-2y}(4280.-17000.y+26400.y^{2}-20900.y^{3}+9250.y^{4}-2370.y^{5}+348.y^{6}-27.8y^{7}+1.05y^{8}-0.0137y^{9}+0.0000284y^{10})$
$F_{\Sigma^{\prime\prime}}^{(p,n)}=e^{-2y}(-0.000933+0.00115y-0.00851y^{2}+0.00811y^{3}-0.00352y^{4}+0.0000644y^{5}+0.00036y^{6}-0.000128y^{7}+0.0000185y^{8}-8.1\times 10^{-7}y^{9}+1.54\times 10^{-11}y^{10})$
$F_{\Sigma^{\prime}}^{(p,n)}=e^{-2y}(-0.00187+0.0274y-0.124y^{2}+0.206y^{3}-0.16y^{4}+0.0672y^{5}-0.0161y^{6}+0.00221y^{7}-0.000159y^{8}+4.77\times 10^{-6}y^{9}+1.01\times 10^{-9}y^{10})$
$F_{\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(176.-490.y+524.y^{2}-276.y^{3}+78.1y^{4}-12.y^{5}+0.935y^{6}-0.0303y^{7}+0.000177y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(p,n)}=e^{-2y}(-0.0278+0.0407y-0.0221y^{2}+0.00217y^{3}+0.00162y^{4}-0.000525y^{5}+0.0000548y^{6}-1.43\times 10^{-6}y^{7}-3.34\times 10^{-8}y^{8})$
$F_{\Delta}^{(p,n)}=e^{-2y}(0.188-0.488y+0.526y^{2}-0.287y^{3}+0.0828y^{4}-0.0124y^{5}+0.000881y^{6}-0.0000216y^{7}+7.34\times 10^{-9}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(p,n)}=e^{-2y}(-1010.+3430.y-4490.y^{2}+2940.y^{3}-1060.y^{4}+215.y^{5}-24.y^{6}+1.36y^{7}-0.0314y^{8}+0.0000709y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(p,n)}=e^{-2y}(0.0706-0.37y+0.625y^{2}-0.512y^{3}+0.225y^{4}-0.0545y^{5}+0.00706y^{6}-0.000454y^{7}+0.0000113y^{8}+2.73\times 10^{-9}y^{9})$
$F_{M\Phi^{\prime\prime}}^{(n,p)}=e^{-2y}(-743.+2510.y-3250.y^{2}+2110.y^{3}-751.y^{4}+150.y^{5}-16.3y^{6}+0.848y^{7}-0.015y^{8}+0.0000709y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.00497+0.0599y-0.158y^{2}+0.164y^{3}-0.0838y^{4}+0.0234y^{5}-0.00359y^{6}+0.00028y^{7}-8.51\times 10^{-6}y^{8}+2.73\times 10^{-9}y^{9}$
$F_{M}^{(n,n)}=e^{-2y}(6050.-26100.y+44400.y^{2}-39100.y^{3}+19600.y^{4}-5850.y^{5}+1040.y^{6}-105.y^{7}+5.58y^{8}-0.125y^{9}+0.000963y^{10})$
$F_{\Sigma^{\prime\prime}}^{(n,n)}=e^{-2y}(4.51\times 10^{-6}-8.31\times 10^{-6}y+0.000803y^{2}-0.00169y^{3}+0.0016y^{4}-0.000798y^{5}+0.000224y^{6}-0.0000347y^{7}+2.67\times 10^{-6}y^{8}-7.76\times 10^{-8}y^{9}+1.46\times 10^{-9}y^{10})$
$F_{\Sigma^{\prime}}^{(n,n)}=e^{-2y}(9.02\times 10^{-6}-0.000196y+0.00242y^{2}-0.00533y^{3}+0.00506y^{4}-0.00254y^{5}+0.000758y^{6}-0.000139y^{7}+0.0000153y^{8}-8.33\times 10^{-7}y^{9}+1.95\times 10^{-8}y^{10})$
$F_{\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(339.-1060.y+1290.y^{2}-788.y^{3}+265.y^{4}-50.4y^{5}+5.32y^{6}-0.286y^{7}+0.00602y^{8})$
$F_{\tilde{\Phi}^{\prime}}^{(n,n)}=e^{-2y}(0.0514-0.123y+0.123y^{2}-0.0637y^{3}+0.0189y^{4}-0.00332y^{5}+0.000335y^{6}-0.0000177y^{7}+3.75\times 10^{-7}y^{8})$
$F_{\Delta}^{(n,n)}=e^{-2y}(0.0129-0.036y+0.0427y^{2}-0.0268y^{3}+0.00939y^{4}-0.0018y^{5}+0.000175y^{6}-7.44\times 10^{-6}y^{7}+1.25\times 10^{-7}y^{8})$
$F_{M\Phi^{\prime\prime}}^{(n,n)}=e^{-2y}(-1430.+5330.y-7740.y^{2}+5720.y^{3}-2370.y^{4}+572.y^{5}-79.4y^{6}+6.y^{7}-0.213y^{8}+0.00241y^{9})$
$F_{\Sigma^{\prime}\Delta}^{(n,n)}=e^{-2y}(-0.000341+0.00418y-0.0111y^{2}+0.0125y^{3}-0.00739y^{4}+0.00247y^{5}-0.000477y^{6}+0.0000503y^{7}-2.49\times 10^{-6}y^{8}+4.92\times 10^{-8}y^{9})$
133Cs Occupation Number Table
The occupation numbers $n^{N}_{nlj}$ of each valence single-nucleon orbital are given in order of increasing orbit, with $N=\{p,n\}$ denoting protons and neutrons respectively, and $n$ begins at $1$. “Max Occ” refers to the maximum occupation number allowed for each orbital. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
SN100PN SN100PN* GCN5082 GCN5082*
$n$ $l$ $j$ Max Occ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$
1 4 7/2 8 3.32 7.73 3.36 7.75 2.90 7.71 2.96 7.70
2 2 5/2 6 1.04 6.00 1.03 5.67 1.38 6.00 1.35 5.73
2 2 3/2 4 0.23 2.63 0.22 2.77 0.30 2.68 0.27 2.77
3 0 1/2 2 0.14 1.61 0.12 1.51 0.14 1.73 0.12 1.64
1 5 11/2 12 0.26 10.02 0.28 10.28 0.28 9.87 0.30 10.14

### v2

@@
 Abstract
-Any discovery of ‘new physics’ in the neutrino sector first requires a precise prediction of the expected Standard Model cross section.
-Currently, Coherent Elastic neutrino-Nucleus Scattering (CEvNS) experiments are statistics limited.
-However, as new and future experiments scale up, it will be necessary to improve the theoretical predictions.
+Any discovery of ‘new physics’ in the neutrino sector will first require a precise baseline prediction of the expected Standard Model cross section.
+While Coherent Elastic neutrino-Nucleus Scattering (CEvNS) experiments are currently statistics limited, upcoming and future experiments at larger scales will be systematics limited.
+At this point it will be necessary to improve the theoretical predictions to improve the experimental sensitivity.
 Here we review the calculation of the CEvNS cross section in a consistent theory of hadronic currents and compute the relevant nuclear form factors using the nuclear shell model.
@@
 We then refine the Standard Model predictions for the recent experimental results of the COHERENT experiment.
-We find that our cross sections are in good agreement with previous predictions, but with significantly smaller uncertainties - by up to a factor of 10.
+We find that our cross sections are in good agreement with previous predictions, but with significantly smaller uncertainties.
 Near-future CEvNS experiments will meaningfully benefit from improved predictions through an increased sensitivity to new-physics signals.
@@
 Steady advancement in detector technology enabled CEvNS to finally be observed in 2017.
-The first detection of neutrinos via CEvNS was performed by the COHERENT experiment using a small (14.6 kg) cesium-iodide detector [1].
-This measurement was quickly followed up by a detection using a liquid argon detector [2].
-The next planned measurement will be with an array of germanium detectors (18 kg) [3].
-All of these measurements were carried out at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory.
-There, neutrinos with energies up to $\sim$50 MeV are created from pions decaying at rest within the mercury spallation target.
-Together, these measurements have observed hundreds of CEvNS events and are broadly consistent with the Standard Model expectation for the process.
+[…]
 CEvNS has many potential applications, both practical and fundamental.
@@
 Notably, CEvNS experiments are the leading probe of some realizations of sterile neutrinos and non-standard neutrino interactions [12].
-CEvNS could be detected in dark matter experiments and potentially provide the first warning of a supernova signal [13].
-Key to maximising the effectiveness of these applications is having precise predictions for the scattering rate.
+CEvNS will also be detected in dark matter experiments, where it is an irreducible background to dark matter searches, giving rise to the so-called neutrino floor [13] or fog [14].
+Recently, XENONnT and PandaX reported the first indication of CEvNS due to 8B solar neutrinos [15, 16].
+Key to maximizing the effectiveness of these applications is having precise predictions for the scattering rate.
 That is, if one is interested in detecting deviations from the Standard Model at the sub-percent level, then a commensurate level of precision is required for the theoretical predictions.
-[…]
-In this work we calculate the CEvNS cross sections for nuclear targets used by COHERENT, including the nuclear structure effects, which are typically ignored.
-To this end, we use large-scale nuclear shell model calculations to derive all relevant nuclear form factors.
-Using several different nuclear Hamiltonians, we evaluate the uncertainty in the cross section from the nuclear structure effects.
-These results are then applied to improve the theoretical predictions of the CEvNS event rates that COHERENT has observed to date.
-Figure 1: Theoretical predictions of SNS-flux averaged cross sections for the three targets considered in this work.
-Shown are our main results using the shell model and Skyrme-Hartree-Fock (SHF) and comparisons with the Klein-Nystrand form factor (using electric charge radii) and COHERENT’s prediction.
-Also shown is COHERENT’s experimental result (black dashed, with gray region showing uncertainty), and a potential future measurement with 5% error (purple hatched).
+The dominant systematic uncertainty in CEvNS rate predictions comes from the neutrino flux, which is presently $\sim 10\%$ for the pion-decay-at-rest neutrinos produced at the SNS.
+However, future dedicated measurements by COHERENT aim to reduce this significantly, to around 2-3% [17].
+With this improvement, cross section uncertainties, currently around 2% [2], will provide a significant contribution to the overall systemic uncertainty in the predicted event rates.
+Relatively little attention has been given to the theoretical uncertainties of the cross section.
+With CEvNS being a low-energy process, it is expected that predictions that treat the nucleus as an elementary particle are sufficient.
+The nuclear structure is then only accounted for through a simple charge form factor.
+Recently, the cross section calculation has been revisited and a modern accounting of nucleon currents applied [18]; however, this formalism has not yet been adopted for use in predicting experimental observables.
+That said, cross section calculations using alternative approaches, using coupled cluster theory [19] (for argon only) and Skyrme-Hartree-Fock theory [20] (for carbon, oxygen, argon, iron and lead) have been performed.
+In this work we present theoretical predictions for the CEvNS cross sections, based on detailed nuclear structure calculations, for targets used by the COHERENT experiment (argon, germanium, cesium and iodine).
+In section II we review the standard calculation of the CEvNS cross section and present two methods to derive the nuclear form factors.
+In section III we revise the full CEvNS cross section calculation (following reference [18]) and present our computations of the relevant nuclear form factors based on the nuclear shell model.
+In section IV we use our form factors to improve the theoretical predictions of the CEvNS flux averaged cross sections and event rates for the COHERENT experiment.
+Finally, in section V we conclude.
 II CEvNS cross section
@@
 With the mass of the neutrinos safely neglected, the maximum recoil an incoming neutrino of energy $E_{\nu}$ can impart is then
-$E_{\nu}^{\rm{max}}=\frac{2E_{\nu}^{2}}{m_{T}+2E_{\nu}}.$ (2)
+$E_{R}^{\rm{max}}=\frac{2E_{\nu}^{2}}{m_{T}+2E_{\nu}}.$ (2)
 The cross section is typically calculated by treating the nucleus as composed of point-like nucleons and accounting for their spatial extent within the nucleus via a form factor.
-This results in a standard form of the CEvNS cross section
+This results in a standard expression for the CEvNS cross section
 $\frac{d\sigma}{dE_{R}}=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{E_{R}m_{T}}{2E_{\nu}^{2}}\right)Q_{\text{w}}^{2}F^{2}(q^{2}),$ (3)
-where $Q_{\text{w}}=Q_{p}Z+Q_{n}(A-Z)$ is the weak nuclear charge in terms of the proton and neutron weak charges and $F(q)$ is the nuclear form factor, typically taken to be of the Helm [20] or Klein-Nystrand [21] form.
+where $Q_{\text{w}}=Q_{\text{w}}^{p}Z+Q_{\text{w}}^{n}(A-Z)$ is the weak nuclear charge in terms of the proton and neutron weak charges and $F(q)$ is the nuclear form factor, typically taken from the Helm [22] or Klein-Nystrand [23] model.
 The above is a good approximation to the CEvNS cross section, however, several improvements are variously made in the literature, which can have percent-level corrections.
@@
 $\frac{d\sigma_{T\nu_{i}}}{dE_{R}}=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{E_{R}m_{T}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)$
-×$\times\left(Q_{p,\nu_{i}}F_{p}(q^{2})+Q_{n}F_{n}(q^{2})\right)^{2}.$ (4)
-Here the flavor dependent charges are taken from [23]:
-$Q_{n}$ $=$ $-1.02352(25)$
-$Q_{p,\nu_{e}}$ $=$ $0.0747(34),\,\,\,\,Q_{p,\nu_{\mu}}=0.0582(34).$ (5)
-Using this equation requires knowledge of the proton and neutron distributions within the nucleus.
-While the former can be measured well due to its electric charge, the latter is not well constrained experimentally.
-Given the much larger weak charge of the neutron, the relatively unknown neutron form factor thus contributes the largest uncertainty to determinations of the CEvNS cross section.
-III Nuclear form factors
+×$\times\left(Q_{\text{w}}^{p,\nu_{i}}F_{p}(q^{2})+Q_{\text{w}}^{n}F_{n}(q^{2})\right)^{2}.$ (4)
+Here we take the flavor dependent charges from [25]:
+$Q_{\text{w}}^{n}$ $=$ $-1.02352(25)$
+$Q_{\text{w}}^{p,\nu_{e}}$ $=$ $0.0747(34),\,\,\,\,Q_{\text{w}}^{p,\nu_{\mu}}=0.0582(34),$ (5)
+and use them in all further calculations.
+Using Eq.
+(II) requires knowledge of both the proton and neutron form factors.
+While the former can be inferred from the electric charge distribution, the latter is not well constrained experimentally.
+Given that the weak charge of the neutron is much larger than the proton’s, we can deduce that the dominant contribution to the uncertainty of the CEvNS cross section comes from the neutron form factor.
+Table 1: Experimental charge rms radii $r_{\text{rms}}^{p}$ values (in fm) for 40Ar, 70,72,73,74,76Ge, 127I and 133Cs, from [28].
+40Ar Ge 127I 133Cs
+70 72 73 74 76
+$r_{\text{rms}}^{p}$ 3.4274(26) 4.0414(12) 4.0576(13) 4.0632(14) 4.0742(12) 4.0811(12) 4.7500(81) 4.8041(46)
+II.1 Nuclear form factors
 Nuclear form factors are defined as the Fourier transform of density distributions of the nucleus:
@@
 electric charge-, weak charge- or mass-density, results in different form factors (and rms radii) which are appropriate for use in different scattering processes.
+When modelling the nucleus as a bound state of point-like nucleons, the neutron and proton point radii, $R_{n,p}$, are obtained from their respective distributions via
+$R_{n}^{2}=\frac{1}{A-Z}\int d^{3}r\,r^{2}\rho_{n}(\mathbf{r}),$ (8)
+and
+$R_{p}^{2}=\frac{1}{Z}\int d^{3}r\,r^{2}\rho_{p}(\mathbf{r}).$ (9)
 Neutral-current neutrino scattering proceeds through the coupling of the Z-boson to the weak charge of the nucleus.
 Therefore CEvNS requires the use of weak-charge form factors which in turn require knowledge of the weak-charge density.
-We note that some past work has used experimentally determined electric-charge rms radii to construct form factors for CEvNS, but this can differ significantly from the weak-charge radii [26].
-It is possible to measure the weak charge distribution through parity-violating electron scattering, which has been performed at Jefferson lab by the PREX and CREX experiments, on lead-208 and calcium-48 respectively.
+It is possible to measure the weak charge distribution directly through parity-violating electron scattering.
+This has been performed at Jefferson lab by the PREX and CREX experiments, on 208Pb and 48Ca targets, respectively.
 However, these experiments have only measured the weak charge at a single momentum transfer: $q=0.76$ fm-1 and 1.28 fm-1, respectively [29, 30].
@@
 the electric charge distributions, or theoretical modelling.
+While electric-charge rms radii are measured to high-precision, they can differ significantly from the weak-charge radii [31] and form factors [20].
+Therefore they cannot be used to produce accurate CEvNS cross sections.
 In this work we will compare several methods for modelling the CEvNS form factors and determine their effect on the resulting cross sections.
 This includes theoretical modelling using both Skyrme-Hartree-Fock (SHF) calculations and the nuclear shell model.
-The SHF results will allow us to accurately model the nucleon distributions, from which rms radii are computed.
+The SHF results will allow us to model the nucleon distributions accurately, from which rms radii are computed.
 The shell model calculations will allow us to include sub-leading nuclear effects through the chiral EFT formalism of [18].
 For comparison with previous work in the literature, we will also perform a data-driven calculation based on electric-charge radii.
-When modelling the nucleus as a bound state of point-like nucleons, the neutron and proton point radius distributions $R_{n},p$ are obtained from
-$R_{n}^{2}=\frac{1}{A-Z}\int d^{3}r\,r^{2}\rho_{n}(\mathbf{r}),$ (8)
-$R_{p}^{2}=\frac{1}{Z}\int d^{3}r\,r^{2}\rho_{p}(\mathbf{r}),$ (9)
-respectively.
-Given that stable nuclei exhibit a common, almost uniform, saturating density it is possible to parameterise their density distributions (to very good accuracy) with fits to distributions of just two parameters: one which defines the size of the nucleus and another which defines the thickness of the ‘skin’ region at the surface.
-At low momentum transfers, the form factor remains relatively insensitive to differences in these parameterizations.
-To fit observation to a density distribution, it can be sufficient to equate a measured or theoretically determined rms radius with that of the parameterized distribution.
-We will make use of this matching process below.
-III.1 From electric-charge densities
-The most reliable access we have to the nuclear size and shape is via electromagnetic probes which couple to the electric charge distribution.
-Protons, being charged, contribute more to this distribution than the neutrons do.
-However, the reverse is true for the weak-charge distribution where neutrons dominate in their contribution.
-This makes the electric charge distribution an imperfect surrogate for the weak-charge distribution.
-To account for this difference, previous works have made some assumptions about the neutron distribution based on the proton distribution.
-In [22] the neutron rms radius was chosen to lie between the proton rms charge radius and an additional 0.3 fm.
-In [29], a fit is made to the proton-neutron radii difference data, however the large scatter in the data used in the fit tends to exaggerate the uncertainty for any one nucleus.
-In this work we will follow the method of [22].
-For our electric-charge inspired distributions we will use the Klein-Nystrand (KN) [21] model which has been previously used in CEvNS research (including by the COHERENT collaboration).
-The Klein-Nystrand form factor is based on both a Yukawa potential of range $a_{k}$ and a hard sphere distribution with radius $r_{A}$.
-It has the form
-$F_{\text{KN}}(q)=3\frac{\sin(qr_{A})-qr_{A}\cos(qr_{A})}{(qr_{A})^{3}}\left[1+(qa_{k})^{2}\right]^{-1},$ (10)
-where $a_{k}=0.7$ fm and $r_{A}$ is the nuclear radius. The nuclear radius is obtained by equating the rms radius radius for the distribution:
-$\langle r^{2}\rangle_{KN}=\frac{3}{5}r_{A}^{2}+6a_{k}^{2}.$ (11)
-with the proton rms charge radius.
-The proton rms charge radii for the target nuclei $\sqrt{\langle r_{p}^{2}\rangle}\equiv r_{\text{rms}}^{p}$ can be obtained from various experimental sources.
-The values we adopt are shown in Tab.
-1.
+II.2 Empirically derived form factor
+The most reliable measurements of nuclear size come from probes of the nuclear electric-charge distribution, e.g., from electron scattering data.
+Protons, being charged, contribute more to the distribution than the neutrons do.
+However, the reverse is true for the weak-charge distribution (where the neutron contribution dominates), making the electric-charge distribution an imperfect surrogate for the weak-charge distribution.
+Therefore, while nuclear charge-radius measurements can help define the proton distribution, additional information (or assumptions) are required to define the neutron distribution.
+While it is difficult to probe the neutron distribution directly, it is possible to measure the proton-neutron radius difference, $\Delta r_{np}=r_{rms}^{n}-r_{rms}^{p}$, via nuclear spectroscopy of antiprotonic atoms [32].
+Presently, these data have large uncertainties, however, an approximately linear relationship between $\Delta r_{np}$ and $I\equiv(N-Z)/A$ is apparent.
+A global fit to the available data is given in reference [33]:
+$\Delta r_{np}=[(0.82\pm 0.54)I-(0.02\pm 0.08)]\,\rm{fm}.$ (10)
+Here we disregard the small uncertainty values given by Eq.
+(10), as we don’t believe the simple linear relationship captures the true scatter that exists in the data.
+However, we can use this formula to estimate that $\Delta r_{np}=0.050\,\rm{fm}$ for 70Ge ($I=0.09$) and $\Delta r_{np}=0.15\,\rm{fm}$ for 136Xe ($I=0.21$).
+All other nuclei in our study lie within this range of $I$ values.
+Therefore, following [24], we conservatively assume that the neutron rms radius lies within the range $r^{p}_{rms}<r^{n}_{rms}<r^{p}_{rms}+0.3$ fm.
+To define our empirical form factors we take $r^{p}_{rms}$ directly from electric-charge radii measurements, and use the range of $r^{n}_{rms}$ to define an uncertainty band.
+For the form factor we use the Klein-Nystrand (KN) [23] model, which has been previously used in CEvNS research (including by the COHERENT collaboration).
+The KN model is based on a hard sphere distribution with radius $r_{A}$ and a Yukawa potential of range $a_{k}$.
+The form factor is given by:
+$F_{\text{KN}}(q)=3\frac{\sin(qr_{A})-qr_{A}\cos(qr_{A})}{(qr_{A})^{3}}\left[1+(qa_{k})^{2}\right]^{-1}.$ (11)
+Here we take $a_{k}=0.7$ fm and obtain $r_{A}$ by equating the rms radius for the distribution with the proton rms charge radius:
+$\langle r_{p}^{2}\rangle\equiv\langle r^{2}\rangle_{KN}=\frac{3}{5}r_{A}^{2}+6a_{k}^{2}.$ (12)
+The proton rms charge radii for the target nuclei are shown in table 1.
 The neutron rms radii are then taken to lie in the range $r_{\text{rms}}^{p}\leq r_{\text{rms}}^{n}\leq r_{\text{rms}}^{p}+0.3$ fm for Ge, I and Cs, whilst for Ar the upper limit is taken to be $r_{\text{rms}}^{p}+0.2$ fm.
-We use this range to bracket the uncertainty in the neutron distribution.
-Table 1: Experimental charge rms radii $r_{\text{rms}}^{p}$ values (in fm) for 40Ar, 70,72,73,74,76Ge, 127I and 133Cs, obtained from [30].
-40Ar Ge 127I 133Cs
-70 72 73 74 76
-$r_{\text{rms}}^{p}$ 3.4274(26) 4.0414(12) 4.0576(13) 4.0632(14) 4.0742(12) 4.0811(12) 4.7500(81) 4.8041(46)
-III.2 From Skyrme-Hartree-Fock distributions
-To model the weak-charge distribution of the nucleus more accurately, we perform Skyrme-Hartree-Fock calculations to find the point-like nucleon distributions within the nucleus.
-While we have experimental probes of the electric-charge distributions, which dominantly probes the protons, for consistency we will also use the theoretical values for $R_{p}$ in this calculation.
+II.3 Skyrme-Hartree-Fock distributions
+To model the weak-charge distribution of the nucleus more directly, we perform Skyrme-Hartree-Fock calculations to find the point-like nucleon distributions within the nucleus.
 The weak radius of the nucleus can then be calculated from:
-$R_{W}^{2}=\frac{1}{Q_{\text{w}}}\left(ZQ_{p}(R_{p}^{2}+r_{\text{w},p}^{2})+(A-Z)Q_{n}(R_{n}^{2}+r_{\text{w},n}^{2})\right)$ (12)
+$R_{W}^{2}=\frac{1}{Q_{\text{w}}}\left[ZQ_{\text{w}}^{p}(R_{p}^{2}+r_{\text{w},p}^{2})+(A-Z)Q_{\text{w}}^{n}(R_{n}^{2}+r_{\text{w},n}^{2})\right]$ (13)
 where $r_{\text{w},p}=1.545(17)$ fm and $r_{\text{w},n}=0.8434(12)$ fm are the weak radii of the proton and neutron, respectively.
-The nucleon densities were evaluated with the skyax solver [31] assuming axial symmetry for several parametrizations of the Skyrme functional, namely SkM, SLy6, SV-min and UNEDF1 as defined in skyax.
+In this equation we have dropped the flavor dependence for visual clarity.
+We take it as understood that all charges and radii are neutrino-flavor dependent and retain that dependence in all calculations.
+The nucleon densities were evaluated with the skyax code [34] assuming axial symmetry for several parametrizations of the Skyrme functional, namely SkM, SLy6, SV-min and UNEDF1 as defined in skyax.
 Pairing correlations are treated at the BCS level with a “density-dependent delta interaction” (DDDI) for SV-min and SkM, and with a “volume delta interaction” (VDI) for SLy6 and UNEDF1.
@@
 The results are summarised in Table 2.
-For simplicity we then parameterize the distribution via the symmetrized 2-parameter Fermi (Fermi) model [32], which has been shown to provide a reasonable fit to similar calculations [26].
+For simplicity we then parameterize the distribution via the symmetrized 2-parameter Fermi model [35], which has been shown to provide a reasonable fit to similar calculations [31].
 The Fermi distribution is defined by:
@@
 $\rho_{0}=\frac{3Q_{\text{w}}}{4\pi c(c^{2}+\pi^{2}a^{2})}.$ (16)
-The (normalised) form factor is then:
+The (normalized) form factor is then:
 $F_{F}(q)$ $=\frac{3}{qc}\left(\frac{\sin(qc)}{(qc)^{2}}\frac{\pi qa}{\tanh(\pi qa)}-\frac{\cos(qc)}{qc}\right)$ (17)
-×$\times\frac{\pi qa}{\sinh(\pi qa)}\frac{1}{1+(\pi a/c)^{2}}$
+×$\times\frac{\pi qa}{\sinh(\pi qa)}\frac{1}{1+(\pi a/c)^{2}}.$
 Motivated by the results for the weak skin of calcium and lead in [31], in this work we set $a=0.55\pm 0.05$ fm.
@@
 (13).
-Table 2: Skyrme-Hartree-Fock predictions of rms point-like proton and neutron radii and the resulting predicted RMS Weak radius.
+Table 2: Skyrme-Hartree-Fock predictions of rms point-like proton and neutron radii and the resulting predicted rms Weak radius (for $\nu_{e}$).
 Nucleus $R_{p}$ (fm) $R_{n}$ (fm) $R_{W}$ (fm)
@@
 133Cs 4.750(12) 4.887(8) 4.957(9)
-IV Detailed nuclear treatment
+III Detailed nuclear treatment
 The formalism discussed thus far has been derived top-down, dealing with the most dominant contributions of neutrino-nucleus couplings.
@@
 This task has been performed for the CEvNS process, in full generality, in [18].
-The resulting cross section is much like the traditional form of Eq.(3), with an additional term for the axial form factor:
+The resulting cross section is similar to Eq.
+(3), with an additional term for the axial form factor:
 $\begin{split}\frac{d\sigma_{A}}{dE_{R}}&=\frac{G_{F}^{2}m_{T}}{4\pi}\left(1-\frac{m_{T}E_{R}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)Q_{\text{w}}^{2}|F_{\text{w}}(\mathbf{q}^{2})|^{2}\\ &+\frac{G_{F}^{2}m_{T}}{4\pi}\left(1+\frac{m_{T}E_{R}}{2E_{\nu}^{2}}-\frac{E_{R}}{E_{\nu}}\right)F_{A}(\mathbf{q}^{2}),\end{split}$ (18)
-However, the full form of the nuclear form factors go beyond the density distributions discussed previously. The Weak form factor takes the form
+However, the full nuclear form factor goes beyond the density distributions discussed previously. The Weak form factor becomes
 $\begin{split}F_{\text{w}}(\mathbf{q}^{2})&=\frac{1}{Q_{\text{w}}}\bigg{[}\bigg{(}Q_{\text{w}}^{p}\Big{(}1+\frac{\langle r_{E}^{2}\rangle^{p}}{6}t+\frac{1}{8m_{N}^{2}}t\Big{)}\\ &+Q_{\text{w}}^{n}\frac{\langle r_{E}^{2}\rangle^{n}+\langle r_{E,s}^{2}\rangle^{N}}{6}t\bigg{)}\mathcal{F}^{M}_{p}(\mathbf{q}^{2})\\ &+\bigg{(}Q_{\text{w}}^{n}\Big{(}1+\frac{\langle r_{E}^{2}\rangle^{p}+\langle r_{E,s}^{2}\rangle^{N}}{6}t+\frac{1}{8m_{N}^{2}}t\Big{)}\\ &+Q_{\text{w}}^{p}\frac{\langle r_{E}^{2}\rangle^{n}}{6}t\bigg{)}\mathcal{F}^{M}_{n}(\mathbf{q}^{2})\\ &-\frac{Q_{\text{w}}^{p}(1+2\kappa^{p})+2Q_{\text{w}}^{n}(\kappa^{n}+\kappa_{s}^{N})}{4m_{N}^{2}}t\mathcal{F}^{\Phi^{\prime\prime}}_{p}(\mathbf{q}^{2})\\ &-\frac{Q_{\text{w}}^{n}(1+2\kappa^{p}+2\kappa_{s}^{N})+2Q_{\text{w}}^{p}\kappa^{n}}{4m_{N}^{2}}t\mathcal{F}^{\Phi^{\prime\prime}}_{n}(\mathbf{q}^{2})\bigg{]},\end{split}$ (19)
 where $m_{N}$ is the nucleon mass and the remaining constants are provided in [18].
-The weak form factor hence depends on the coherent and semi-coherent responses $F^{M}_{N}$ and $F^{\Phi^{\prime\prime}}_{N}$, respectively, where $N=\{p,n\}$ denotes protons and neutrons.
-The axial-vector form factor has the form
+The weak form factor hence depends on the coherent and semi-coherent responses $\mathcal{F}^{M}_{N}$ and $\mathcal{F}^{\Phi^{\prime\prime}}_{N}$, respectively, where $N=\{p,n\}$ denotes protons and neutrons.
+The axial form factor is given by
 $\begin{split}F_{A}&(\mathbf{q}^{2})=\frac{8\pi}{2J+1}\\ \times&\left((g_{A}^{s,N})^{2}S_{00}^{\mathcal{T}}(\mathbf{q}^{2})-g_{A}g_{A}^{s,N}S_{01}^{\mathcal{T}}(\mathbf{q}^{2})+(g_{A})^{2}S_{11}^{\mathcal{T}}(\mathbf{q}^{2})\right)\end{split}$ (20)
@@
 $\begin{split}S_{00}^{\mathcal{T}}&=\sum_{L}\left[\mathcal{F}_{+}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\right]^{2},\\ S_{11}^{\mathcal{T}}&=\sum_{L}\left[[1+\delta^{\prime}(\mathbf{q}^{2})]\mathcal{F}_{-}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\right]^{2},\\ S_{01}^{\mathcal{T}}&=\sum_{L}2\left[1+\delta^{\prime}(\mathbf{q}^{2})\right]\mathcal{F}_{+}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2})\mathcal{F}_{-}^{\Sigma^{\prime}_{L}}(\mathbf{q}^{2}),\end{split}$ (21)
-The correction to the leading SD coupling for the transverse SD response $\delta^{\prime}(\mathbf{q}^{2})$ is given in [15].
-In this more-detailed treatment of the nucleus, the full Weak radius receives additional contributions from spin-orbit effects and the strange electric-charge distribution.
-It takes the form
+The correction to the leading spin-dependent (SD) coupling for the transverse SD response $\delta^{\prime}(\mathbf{q}^{2})$ is given in [18].
+In this more-detailed treatment of the nucleus, the full weak radius receives additional contributions from spin-orbit effects and the strange electric-charge distribution:
 $\begin{split}R_{\text{w}}^{2}&=\frac{ZQ_{\text{w}}^{p}}{Q_{\text{w}}}\bigg{(}R_{p}^{2}+\langle r_{E}^{2}\rangle^{p}+\frac{Q_{\text{w}}^{n}}{Q_{\text{w}}^{p}}\left(\langle r_{E}^{2}\rangle^{n}+\langle r_{E,s}^{2}\rangle^{N}\right)\bigg{)}\\ &+\frac{NQ_{\text{w}}^{n}}{Q_{\text{w}}}\bigg{(}R_{n}^{2}+\langle r_{E}^{2}\rangle^{p}+\langle r_{E,s}^{2}\rangle^{N}+\frac{Q_{\text{w}}^{p}}{Q_{\text{w}}^{n}}\langle r_{E}^{2}\rangle^{n}\bigg{)}\\ &+\frac{3}{4m_{N}^{2}}+\langle\tilde{r}^{2}\rangle_{\text{so}},\end{split}$ (22)
@@
 $\begin{split}\langle\tilde{r}^{2}\rangle_{\text{so}}&=-\frac{3Q_{\text{w}}^{p}}{2m_{N}^{2}Q_{\text{w}}}\Big{(}1+2\kappa^{p}+2\frac{Q_{\text{w}}^{n}}{Q_{\text{w}}^{p}}(\kappa^{n}+\kappa_{s}^{N})\Big{)}\mathcal{F}^{\Phi^{\prime\prime}}_{p}(0)\\ &-\frac{3Q_{\text{w}}^{n}}{2m_{N}^{2}Q_{\text{w}}}\Big{(}1+2\kappa^{p}+2\kappa_{s}^{N}+2\frac{Q_{\text{w}}^{p}}{Q_{\text{w}}^{n}}\kappa^{n}\Big{)}\mathcal{F}^{\Phi^{\prime\prime}}_{n}(0).\end{split}$ (23)
-To apply this formalism and calculate cross sections for a given target we will employ the nuclear shell model to compute the nuclear responses above.
-IV.1 Evaluating the nuclear responses
+To make use of this formalism we need to calculate the nuclear responses $\mathcal{F}_{N}^{M}$, $\mathcal{F}_{N}^{\Phi^{\prime\prime}}$, and $\mathcal{F}_{N}^{\Sigma^{\prime}}$ for our chosen targets – for this we employ the nuclear shell model.
+III.1 Evaluating the nuclear responses
 Figure 1: Spin-independent weak form factors for CEvNS in the four targets relevant to current experiments.
-Bands show the error due to uncertainties in the neutron distribution (Charge) and from using different interactions (SHF and Shell model).
-To compute the weak and axial-vector nuclear form factors in Eqs.
-(18) and (19), we need to evaluate the $\mathcal{F}^{M}_{N}$, $\mathcal{F}^{\Phi^{\prime\prime}}_{N}$ and $\mathcal{F}^{\Sigma^{\prime}}_{N}$ nuclear responses.
-The nuclear responses originate from a multipole decomposition of the lepton-nucleus amplitude, developed in [33, 34].
-Here, they are defined as:
+Bands show the error due to uncertainties in the neutron distribution (Charge) and from using different nuclear structure models (SHF and Shell model).
+The arrows indicate the COHERENT regions of interest (ROI).
+The nuclear responses originate from a multipole decomposition of the lepton-nucleus amplitude, developed in [36, 37]. Here, they are defined as:
 $|\mathcal{F}^{X}_{N}(q^{2})|^{2}\equiv$ $\frac{4\pi}{2J_{i}+1}\sum\limits_{J=0}\langle J_{i}||X_{J}^{(N)}||J_{i}\rangle\langle J_{i}||X^{(N)}_{J}||J_{i}\rangle,$
 $|\mathcal{F}^{\Sigma^{\prime}}_{N}(q^{2})|^{2}\equiv$ $\sum\limits_{J=1}\langle J_{i}||\Sigma_{J}^{\prime(N)}||J_{i}\rangle\langle J_{i}||\Sigma^{\prime(N)}_{J}||J_{i}\rangle,$
-where $J_{i}$ is the nuclear ground state angular momentum, and the nuclear operators $X=M,\,\,\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are given in Appendix C.
+where $J_{i}$ is the nuclear ground state angular momentum, and $X=\{M,\,\,\Phi^{\prime\prime}\}$ and $\Sigma^{\prime}$ are nuclear operators whose full form is given in Appendix C.
 The nature of these operators can be understood from their long-wavelength ($q\rightarrow 0$) limit (provided in Table 3).
@@
 The $\Sigma^{\prime}$ operator is spin-dependent (SD), picking out the transverse projection of the spin current.
-All nuclei have non-zero responses to the $M$ and $\Phi^{\prime\prime}$ operators, while only odd-nuclei with a $J_{i}\neq 0$ ground state have a $\Sigma^{\prime}$ response.
+All nuclei have non-zero responses to the $M$ and $\Phi^{\prime\prime}$ operators, while only odd-nuclei with a $J_{i}\neq 0$ ground state have a non-zero $\Sigma^{\prime}$ response.
 Table 3: Leading order terms of the nuclear operators in the long-wavelength limit $q\rightarrow 0$ [38].
@@
 $\frac{q}{m_{N}}\Phi^{\prime\prime}_{2M}(q\vec{x}_{m})$ $-\frac{q}{m_{N}}\frac{1}{\sqrt{30\pi}}\left[x_{m}\otimes\left(\vec{\sigma}(m)\times\frac{\vec{\nabla}}{i}\right)_{1}\right]_{2M}$
-Using the Wigner-Eckhart theorem the nuclear matrix elements in Eq. (IV.1) can be reduced in angular momentum and isospin:
+Using the Wigner-Eckhart theorem the nuclear matrix elements in Eq. (III.1) can be reduced in angular momentum and isospin, giving:
 $\langle J_{i};TM_{T}\ ||\sum_{m=1}^{A}\hat{O}_{J,\tau}(q\vec{x}_{m})||\ J_{i};TM_{T}\rangle=(-1)^{T-M_{T}}$
@@
 $\begin{split}\Psi^{J;\tau}_{|\alpha|,|\beta|}&\equiv\frac{\langle J_{i};T\ \vdots\vdots\left[a^{\dagger}_{|\alpha|}\otimes\tilde{a}_{|\beta|}\right]_{J;\tau}\vdots\vdots\ J_{i};T\rangle}{\sqrt{(2J+1)(2\tau+1)}},\end{split}$ (27)
-where $\tilde{a}_{\beta}=(-1)^{j_{\beta}-m_{j_{\beta}}+1/2-m_{t_{\beta}}}\ a_{|\beta|;-m_{j_{\beta}},-m_{t_{\beta}}}$ and $\otimes$ denotes a tensor product.
-This basis gives response functions of the form $e^{-2y}p(y)$, where $p(y)$ is a polynomial with $y=(qb/2)^{2}$ and $b={1}/{\sqrt{m_{N}\omega}}\approx\sqrt{41.467/(45A^{-1/3}-25A^{-2/3})}$ fm is the harmonic oscillator length parameter ($\omega$ is the oscillator frequency).
-IV.2 Form factors from shell model calculations
-Large-scale nuclear shell model calculations were performed using NuShellX [37], which is often employed to obtain theoretical energies, wave functions and spectroscopic overlaps for a wide range of nuclei.The program’s inputs include the relevant valence space for the nucleus/isotope of interest - which constitutes a set of single nucleon levels above a completely filled nuclear core - as well as its corresponding list of two-body interactions.
-Each interaction is fitted to a set of experimental nuclear data for a range of nuclei in the relevant valence space.
+where $\tilde{a}_{|\beta|}=(-1)^{j_{\beta}-m_{j_{\beta}}+1/2-m_{t_{\beta}}}\ a_{|\beta|;-m_{j_{\beta}},-m_{t_{\beta}}}$ and $\otimes$ denotes a tensor product.
+This basis gives response functions of the form $e^{-2y}p(y)$, where $p(y)$ is a polynomial of the dimensionless parameter $y=(qb/2)^{2}$ and $b={1}/{\sqrt{m_{N}\omega}}\approx\sqrt{41.467/(45A^{-1/3}-25A^{-2/3})}$ fm is the harmonic oscillator length parameter with $\omega$ as the oscillator frequency.
+Figure 2: Spin-dependent form factors relevant to CEvNS for the three $J_{i}\neq 0$ targets we consider. The curves denote the different shell model interactions as described in Appendix B.
+III.2 Form factors from shell model calculations
+Large-scale nuclear shell model calculations were performed using NuShellX [40], which can produce level energies, wave functions and spectroscopic overlaps for a wide range of nuclei.
+The program’s inputs include the relevant valence space for the nucleus of interest – which constitutes a set of nucleon orbits/levels above a completely filled nuclear core – as well as its corresponding list of two-body interactions.
+Each set of two-body interactions, also referred to as the ‘interaction’, is fitted to selected experimental nuclear data.
 The contemporary approach is to derive the interactions from the nucleon-nucleon interaction using effective field theory.
-The interactions are then fine tuned by fitting a limited number of parameters to a body of nuclear data - usually excitation energies in a range of nuclei in the valence space.
+The interactions are then fine tuned by fitting a limited number of parameters to a body of nuclear data – usually excitation energies in a range of nuclei relevant to the valence space.
 The aforementioned OBDMEs are extracted for the valence space levels, with the values dependent on the particular shell model interactions used.
 In this work we explore a variety of shell model interactions and truncations for our chosen targets: 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs.
-Using the shell model calculations detailed below, we extract the OBDME for all targets.
-The OBDMEs are then used to compute the nuclear responses using the Mathematica package described in [38].
-Analytic form factor expressions for 40Ar, 70,72,73,74,76Ge and 127I are provided in [39], while those for 133Cs are provided in Appendix D.
+Details of the shell model calculations, as well as comparisons with experimental nuclear structure data can be found in Ref.
+[41] and in Appendix B.
+The resulting OBDMEs are then used to compute the nuclear responses with the Mathematica package described in [42].
 The form factors for each of the nuclear responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are plotted as functions of $q$ in Appendix D.
-The form factors are used to derive point-proton, point-neutron and weak radii for all isotopes for comparison with those obtained with our SHF calculations, and are given in Table 7 of Appendix A.
-Additionally, we make comparisons with experimental energy spectra, moments and transitions in Appendix B.
+Analytic expressions of the form factors for 40Ar, 70,72,73,74,76Ge and 127I are provided in [41], while those for 133Cs are provided in Appendix D.
+For comparison with our SHF calculations, the form factors are used to derive point-proton, point-neutron and weak radii for all isotopes, and are given in Table 7 of Appendix A.
 The shell-model derived weak form factors are shown in Fig.
-2, compared to form factors based on the charge distribution and SHF calculations.
+1, compared to the form factors based on the charge distribution and SHF calculations.
 The uncertainty band denotes the range of values obtained given the different shell model interactions.
 The shell model results for each isotope show good agreement with each other (low overall uncertainty) and fall within the large uncertainty band of the form factors derived from the charge distribution.
-They are also generally in very good agreement with the SHF results within the experimentally accessible regions of interest.
-The axial form factor is shown in Fig.
-3 for each interaction individually.
-The sections below outline the valence space, valence space truncation and shell model interactions used in this work for each target.
-Figure 3: Spin-dependent form factors relevant to CEvNS for the three $J_{i}\neq 0$ targets we consider. The curves denote the different shell model interactions as described in the text.
-40Ar
-The $sdpf$ valence space was employed for this nucleus, which constitutes single particle levels $1d_{5/2}$, $2s_{1/2}$, $1d_{3/2}$, $1f_{7/2}$, $2p_{3/2}$, $1f_{5/2}$ and $2p_{1/2}$.
-The valence space truncation employed consisted of an unrestricted $sd$ shell and an empty $pf$ shell for the protons, and a full $sd$ shell with an unrestricted $pf$ shell for the neutrons.
-Four shell model interactions were used, namely SDPF-NR [40], SDPF-U [41], EPQQM [42] and SDPF-MU [43], all of which consider a 16O core.
-This isotope was not considered in the work of [35, 38].
-The 40Ar theoretical energy levels for the above shell model interactions are plotted against the experimental values in Fig.
-5.
-The first two theoretical energy levels for all interactions are consistent with experiment within $\sim 200$ keV, except for SDPF-MU which is within $\sim 400$ keV from experiment for the $2_{1}^{+}$ state.
-Overall, the energy spectrum of the EPQQM interaction is least consistent with the other shell model calculations and with experiment, with the spectrum being much more sparse and spread out in the level spacings.
-The magnetic dipole and electric quadrupole moments and transitions are presented in Table 8, where the effective $g$ factors and charges employed take on values of $e_{p}=1.5$, $e_{n}=0.5$, $g_{sp}=5.586$, $g_{sn}=-3.826$, $g_{lp}=1$, and $g_{ln}=0$.
-Information on the ground state is most important for comparison with theoretical values, as it is used in the scattering calculations here.
-However, as the ground state of 40Ar is $0^{+}$, the magnetic dipole and electric quadrupole moment values are identically zero.
-Nevertheless, we employ the next non-zero state ($2_{1}^{+}$), which can give an indication of the overall accuracy of the calculation through information on the spectrum beyond the ground state.
-It is found that the shell model moments are not consistent with experimental values, and most experimental transitions are poorly reproduced as well.
-The reason for the failure of the shell model for the case of 40Ar is understood: the assumption that 40Ca is an inert core is not justified.
-Its ground state is near spherical, as assumed in the shell model, but it has low-excitation deformed states formed by exciting pairs of protons and neutrons across the shell gaps at $N=20$ and $Z=20$.
-This leads to mixing between the spherical shell model states and the deformed states in nuclei like 40Ar.
-The case of 42Ca, an isotone of 40Ar, is best studied, as has recently been reviewed by Stuchbery and Wood [44].
-They infer that the first 2+ state of 42Ca is about 50% due to the spherical shell model configuration.
-The ground state is about 65% of the shell model configuration.
-Similar mixing is present in 40Ar.
-Thus the shell model calculations performed here are limited by the presence of configuration mixing well beyond the usual valence space.
-The obtained form factors are compared against those based on the SHF calculations and the charge distribution in Appendix D.
-The maximum difference between the form factors for the $M$ response at the minimum $q$ value is $\sim 0.3\%$, and $\sim 11\%$ at the maximum $q$ value.
-For $\Phi^{\prime\prime}$, these values are $\sim 24\%$ and $\sim 21\%$, respectively, where the largest differences occur between the SDPF-MU and EPQQM interactions.
-70,72,73,74,76Ge
-The germanium shell model calculations were performed in the $f_{5}pg_{9}$ model space, which consists of the single-particle levels $2p_{3/2}$, $1f_{5/2}$, $2p_{1/2}$ and $1g_{9/2}$.
-The work of [35, 38] employed the GCN2850 interaction [45], with a valence space truncation that consisted of limiting the occupation number of the $1g_{9/2}$ level to no more than two nucleons above the minimum occupation for all isotopes.
-Here, we consider an unrestricted $f_{5}pg_{9}$ model space with the JUN45 [46] and jj44b [47] interactions.
-The experimental energy spectra are plotted against the theoretical shell model ones for the even isotopes in Fig.
-6 (Appendix B.2).
-The spectra comparison for 73Ge is provided in [39], alongside a brief summary of the two aforementioned shell model interactions, where the theoretical spectra were found to be inconsistent with another, and not particularly consistent with experiment.
-In the cases of 70,72,74Ge, the low-lying experimental spectra are a little better reproduced by the JUN45 interaction results, which was fitted to the Ge isotopes in consideration (the jj44b interaction on the other hand was not fitted to these).
-The shell model spectra for 76Ge are practically identical to one another up to $\sim 1.9$ MeV.
-An agreement between theoretical and experimental levels of $\sim 200$ keV would be considered satisfactory for such shell model calculations using heavier nuclei, however this agreement exists only for some levels of the even isotopes.
-The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided for the even isotopes in Tables 9 and 10.
-Much like with 40Ar, the magnetic and electric moment information is not available for the ground state $0^{+}$, however the subsequent non-zero states are used here as a test of the overall accuracy of the shell model code.
-As before, the values for 73Ge are provided in [39].
-For the magnetic dipole moments of the even isotopes, the experimental $2_{1}^{+}$ and $4_{1}^{+}$ states are well reproduced by both shell model interactions, whereas the jj44b value of the $2_{2}^{+}$ state is consistently closer to experiment compared to JUN45.
-The $2_{1}^{+}$ electric quadrupole moment is quite inconsistent with the experimental values for 70,72,74Ge, where only the jj44b 76Ge value lies within the experimental range.
-The individual germanium form factors are plotted (for the $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ responses) in Appendix D, and compared against the SHF and charge distribution form factors.
-The $M$ response is consistent amongst all calculations, with a maximum difference of $\sim 0.2\%$ at the lower end of the COHERENT $q$ region-of-interest (ROI) for the even isotopes, and $\sim 3\%$ at the higher end for 70,72Ge.
-For 73Ge these take on the values of $\sim 0.3\%$, and $\sim 20\%$ at the higher end for 73,74,76Ge.
-For the $\Phi^{\prime\prime}$ channel, the maximum difference is observed between the GCN2850 and jj44b shell model calculations, which is $\sim 80\%$ for 70Ge at both ends of the $q$ spectrum, factor of $\sim 2.4$ for 72Ge, factor of $\sim 2.3$ for 73Ge, factor of $\sim 2.2$ for 74Ge, and $\sim 60\%$ for 76Ge.
-The difference in the SD $\Sigma^{\prime}$ channel for 73Ge is $\sim 30\%$ at both ends of the $q$ ROI.
-127I
-The iodine calculations were performed in the model space which includes all proton and neutron orbits in the major shell between magic numbers 50 and 82, with single particle levels $1g_{7/2}$, $2d_{5/2}$, $2d_{3/2}$, $3s_{1/2}$ and $1h_{11/2}$.
-The work of [35, 38] utilises a nuclear shell model interaction developed by Baldridge and Dalton [48], which we refer to as “B&D” [39].
-The valence space restriction employed in [35, 38] involves fixing the occupation number of $1h_{11/2}$ to the minimum allowed nucleon number.
-Here, we consider the SN100PN [49] and GCN5082 [45] interactions, and perform two calculations for each.
-Our shell model calculations are performed in the same model space as the B&D calculations of [35, 38], but with different restrictions.
-The model space restrictions employed in our work are the same as those in [39], with the four different calculations given the names “SN100PN” and “GCN5082” (which share the same truncation), as well as “Truncated SN100PN” (with a more truncated scheme), and “GCN5082 Expanded” (with an alternative truncation to “GCN5082”).
-The theoretical energy spectra are plotted against the experimental ones in [39], in addition to the theoretical and experimental electric quadrupole and magnetic dipole moments, with the proton and neutron effective charges and $g$ factors also provided in [39].
-The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and charge distribution form factors in Appendix D.
-The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 30\%$ at the maximum $q$ value.
-For $\Phi^{\prime\prime}$, these values are $\sim 55\%$ and $\sim 50\%$, respectively.
-For $\Sigma^{\prime}$, the factor differences are $\sim 44$ and $\sim 81$, due to the large difference between the Truncated SN100PN calculation and the remaining calculations.
-This is consistent with the integrated form factor (IFF) results in [39], which showed a more pronounced difference between the SD Truncated SN100PN IFF results and other calculations.
-This indicates that a more severe valence space truncation for a shell model interaction has a larger effect compared to a second more comparable truncation.
-133Cs
-Here, we consider the same valence space and shell model interactions (SN100PN and GCN5082) as 127I above.
-We have performed four calculations for 133Cs, two of which employ the same truncations as the “SN100PN” and “GCN5082” calculations in Section IV (and are given the same names here).
-We consider two additional calculations, one each for the SN100PN and GCN5082 interactions, where the neutron $2d_{5/2}$ level now has a minimum of 4 neutrons (as opposed to previously being completely filled), with the remainder of the proton and neutrons levels retaining the previous truncations.
-This is to explore the effect of a less truncated valence space on the final results.
-These two additional calculations are termed “SN100PN Expanded” and “GCN5082 Expanded”.
-The experimental energy spectra are plotted against the theoretical values for all the above calculations in Fig.
-7.
-Only the SN100PN result reproduces the correct ground state spin and parity of $7/2^{+}$, however all four calculations reproduce the correct spins and parities of the three lowest-lying states within $200$ keV of experiment.
-For both shell model interactions, using a less truncated valence space causes the spacing of the levels to widen compared to the original calculations, generally pushing the spectrum further up.
-This is likely due to the opening up of the $2d_{5/2}$ level in these calculations.
-The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided in Tables 11 and 12.
-The proton and neutron effective charges and $g$ factors employed here are the same as those used in [39] for 127I.
-The experimental magnetic dipole moments are much better reproduced by the shell model calculations compared to the electric quadrupole ones.
-The experimental electric quadrupole transitions are only reproduced within the uncertainty margins by the GCN5082 calculation, specifically for transitions from the $5/2_{1}^{+}$ and $3/2_{1}^{+}$ states to the ground state.
-On the other hand, the theoretical magnetic dipole transition values are not consistent with experiment.
-However, the important test of the shell model wavefunction is the magnetic dipole moment, which is sensitive to the wavefunction of the single state, and to the single nucleon structure of the state, rather than to emerging collective properties, which affect the E2 transition rates and the quadrupole moments.
-The magnetic dipole transitions are suppressed among the dominant low-excitation configurations, as is evident from their small values, and so become very sensitive to small component in the wavefunction that are not of interest here.
-Despite the lack of agreement for some observables, the present shell model calculations can be trusted to include the important single-particle configurations in the ground-state wavefunction.
-The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and KN ones in Appendix D.
-The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 20\%$ at the maximum $q$ value.
-For $\Phi^{\prime\prime}$, these values are $\sim 6\%$ and $\sim 7\%$, respectively.
-For $\Sigma^{\prime}$, these differences are $\sim 13\%$ and $\sim 18\%$.
-Table 4: Our CEvNS cross section predictions in units of 10-40cm2
+They are also generally in very good agreement with the SHF form factors within the experimentally relevant regions of interest (i.e.
+the momentum transfer, $q$, which corresponds to observable nuclear recoils in the COHERENT detectors).
+The axial form factors for each target and interaction are shown in Fig.
+2.
+The germanium and cesium axial form factors exhibit good agreement across the different interactions, while iodine has more disparate results.
+However, inspecting Eq.
+(18) and Figs.
+1 and 2, we see that the axial term is a factor of (at least) $10^{3}$ smaller than the spin independent term, and is thus a subdominant contribution.
+This leads us to conclude that the axial form factor uncertainty will be a negligible contribution to the total uncertainty.
+Table 4: Our CEvNS cross section predictions, compared with COHERENT’s predictions and measurements (in units of 10-40cm2)
 COHERENT
@@
 CsI 189(4) 189.3(5) 186.3(5) 189(6) 165(30)
-V Experimental observables
-To date, CEvNS has been observed by the COHERENT collaboration using the stopped pion source of neutrinos from the SNS, with both cesium-iodide [50] and argon [2] detector targets.
-Presently COHERENT have also deployed a germanium detector and measurements are ongoing.
-There are also experiments aiming to measure CEvNS from reactor neutrinos, including CONUS [51], RICOCHET [52] and NUCLEUS [53].
-However, the momentum transfer from reactor neutrinos is very small and thus the form factor has little effect on the cross section.
-We therefore choose to focus our present analysis on the cesium-iodide, argon and germanium targets of COHERENT and leave reactor based experiments to a future work.
-V.1 Flux averaged cross sections
-First we calculate the flux averaged cross sections, by performing a weighted average of the cross sections over the three incoming neutrino flavor spectra and the isotopes in the target:
+IV Experimental observables
+[…]
+Figure 3: Theoretical predictions of SNS-flux averaged cross sections for the three targets considered in this work.
+Shown are our main results using the shell model and Skyrme-Hartree-Fock (SHF) and comparisons with the Klein-Nystrand form factor (using electric charge radii) and COHERENT’s prediction.
+Also shown is COHERENT’s experimental result (black dashed, with gray region showing uncertainty), and a potential future measurement with 5% error (purple hatched).
+IV.1 Flux-averaged cross sections
+First we calculate the flux-averaged cross sections, by performing a weighted average of the cross sections over the three incoming neutrino flavor spectra and the isotopes in the target:
 […]
 Here the neutrino spectra, $\frac{d\phi}{dE_{\nu}}$, are normalized to unity.
-For this calculation we use the neutrino spectra provided in [54], which were derived from detailed GEANT4 simulations of the proton beam colliding in the mercury spallation target.
-The uncertainties in the cross sections are determined from two sources: the form factors and the nucleon charges.
-We neglect uncertainties in the modelling of the neutrino spectra, which are unavailable, but note that errors in the spectra should affect all of the compared cross sections approximately the same.
-The form factor uncertainties are found by computing the flux averaged cross section with our two extremal form factors for each model (i.e.
-the upper and lower limits of the bands shown in Fig.
-2).
-The central value and uncertainty is then taken as ($\sigma_{\rm{max}}+\sigma_{\rm{min}}$)/2 and ($\sigma_{\rm{max}}-\sigma_{\rm{min}}$)/2, respectively.
-The uncertainty in the nucleon charges is taken from [23].
-The total uncertainty is determined by summing these two sources in quadrature.
+For this calculation we use the neutrino spectra provided in reference [47], which were derived from detailed GEANT4 simulations of the proton beam colliding in the mercury spallation target.
 The KN and SHF models are calculated using Eq.
@@
 (18).
-The results of this calculation are shown in Fig.
-1 and provided numerically in table 4.
-Our argon results are all in very close agreement with each other, owing to the fact that argon is an even-even nucleus which doesn’t receive large spin-orbit, and no spin-dependent, contributions.
-This result is slightly larger than the COHERENT prediction, though with the large uncertainties, is still in reasonable agreement.
-Our detailed nuclear results do shrink the uncertainty band, though it will be some time before the statistical errors in the experimental results make this difference relevant.
-Our cesium-iodide results are in good agreement with the COHERENT prediction.
-Here we do see a small difference between our shell model results and the SHF calculations.
-While both cesium and iodine are odd nuclei, the axial term is a positive contribution and so the resulting difference stems from the spin-orbit term.
-Our germanium results show a similar trend to the cesium-iodide results, with the shell model results being slightly smaller than the SHF results, which are both with the prior uncertainty derived from charge distributions.
-Presently there are no COHERENT results to compare with, though they are expected in the near future.
+The uncertainties in our cross section calculations are determined from two sources: the form factors and the nucleon charges.
+We neglect uncertainties in the modeling of the neutrino spectra, which are unavailable, but note that errors in the spectra should affect all of the compared cross sections approximately the same.
+The form factor uncertainties are found by computing the flux averaged cross section with our two extremal form factors for each model (i.e.
+the upper and lower limits of the bands shown in Fig.
+1).
+The central value and uncertainty are then taken as ($\sigma_{\rm{max}}+\sigma_{\rm{min}}$)/2 and ($\sigma_{\rm{max}}-\sigma_{\rm{min}}$)/2, respectively.
+The uncertainty in the nuclear weak charge is propagated from the values for the nucleon charges given in Eq.
+(5) (taken from [25]).
+The total uncertainty is determined by summing the two sources in quadrature.
+We find that for our shell model and SHF calculations, the form factor uncertainties are sub-dominant to the nuclear charge uncertainties, with the latter contributing $\sim 0.3\%$.
+Figure 3 shows the main result of this work: the flux-averaged cross sections for argon, germanium, and cesium-iodide, calculated using our three approaches.
+These results are compared with COHERENT’s theoretical prediction (in black) and experimental result (black dashed, with gray area showing the uncertainty band), from [43, 2].
+In the case of germanium, our prediction can be used to compare with future COHERENT results, when they become available.
+The results are also provided numerically in table 4.
+The argon cross sections are all in very close agreement with each other, and are slightly larger than the COHERENT prediction, but given the uncertainties, is still in reasonable agreement.
+Our SHF and shell model results have a much smaller uncertainty band than the COHERENT prediction and their concordance suggests the result is also accurate.
+The cesium-iodide cross sections are also in good agreement with the COHERENT prediction, falling within the uncertainty band, but again, ours have much smaller uncertainties.
+For germanium and cesium-iodide the shell model cross section is smaller than the SHF cross section, with both lying within the uncertainty band of the charge-derived result.
+There are several potential sources for this difference.
+Firstly, cesium-133, iodine-127, and germanium-73 are odd nuclei and thus the cross section receives a contribution from the axial current.
+However, in the previous section we found that the axial nuclear form factor, $F_{A}$, is small.
+Secondly, our shell model result includes the nuclear spin-orbit response, $\mathcal{F}^{\Phi^{\prime\prime}}$, however this enters the cross section $\propto q^{2}$, and thus, is also small.
+To investigate the size of the contributions from the additional nuclear responses, we computed the cross sections using only the shell model derived $F_{M}$ nuclear form factors.
+For argon, cesium-iodide and germanium, we found that including the axial and spin-orbit terms reduces the flux-averaged cross section by $0.05\%$, $0.05\%$ and $0.08\%$, respectively, well below our margin of error.
+This shows that the additional nuclear responses have very little effect on the cross section and can thus be neglected at the $<0.1\%$ level of precision.
+We therefore conclude that the difference between the SHF and shell model cross sections is due to the $F_{M}$ response.
+This can be seen visually in Fig.
+1, where the argon form factor is the same, while the germanium and iodine (but not cesium) shell model form factors decrease faster than the SHF ones.
+Note, while Fig.
+1 only displays the result for germanium-73, the results are very similar for the other isotopes.
 Figure 4: Total event count predictions for the COHERENT experiment’s argon and cesium-iodide detectors, using our 3 cross section calculations.
@@
 […]
-We note that the assumed cross section model will affect the experimentally determined flux averaged cross section through the calculation of the experiment’s sensitivity.
-Thus as measurements become better, one must be cautious to not interpret deviations of experimental results from predictions which used a different cross section model.
-For a direct comparison with the observable experimental results we also compute the predicted number of events for COHERENT’s detectors.
-V.2 Experimental event counts
+We note that one must be cautious when interpreting deviations of experimental results from predictions.
+The experimentally determined flux-averaged cross section is typically calculated assuming a theoretical form of the cross section (e.g.
+through the calculation of an average efficiency factor).
+Where the energy dependence of the assumed cross section differs from the one used to produce a prediction, a fair comparison cannot be made.
+While this effect is likely small, for a direct comparison with the experimental results, we now compute the predicted number of events for the given experimental runs.
+IV.2 Experimental event counts
 The differential scattering rate of neutrinos from nuclei, per target and incoming neutrino, is found by integrating over the incoming neutrino spectra and summing them:
@@
 $E_{\nu}^{\mathrm{min}}=\sqrt{\frac{E_{R}m_{T}}{2}}.$ (30)
-To go from the raw scattering rate to the expected number of counts in an experiment we follow the recipes provided by the COHERENT collaboration in [54] and [55].
+To go from the raw scattering rate to the expected number of counts in an experiment we follow the recipes provided by the COHERENT collaboration in reference [47] and [48].
 The total number of detected CEvNS events is found by integrating over the detected energy taking into account the quenching of nuclear recoils, detector resolution effects and the energy dependent detection efficiency.
-For the cesium-iodide detector this has the form
+For the cesium-iodide detector this is:
 […]
 ×$\times\int dE_{R}P(P\!E|E_{R})\frac{dR}{dE_{\nu}},$
-while for the argon detector the form is
+while for the argon detector it is:
 […]
-×$\times\int dE_{R}\,P(E_{ee}|E_{R})\frac{dR}{dE_{\nu}},$
-where $P\!E$ is the detected photoelectrons, $N_{T}$ is the number of targets in the detector and $\Phi=\frac{\rm{POT}f_{\pi}}{4\pi d^{2}}$ is total number of neutrinos (per flavor) expected at the target [calculated from the total number of protons on target (POT) and the average number of pions produced per POT, $f_{\pi}$].
+×$\times\int dE_{R}\,P(E_{ee}|E_{R})\frac{dR}{dE_{\nu}}.$
+Here $P\!E$ is the detected photoelectrons, $N_{T}$ is the number of targets in the detector and $\Phi$ is total number of neutrinos, per flavor, expected at the detector.
+The latter is computed as $\Phi=\frac{{\rm{POT}}f_{\pi}}{4\pi d^{2}}$, where POT is the total number of protons on target, $f_{\pi}$ is the average number of pions produced per POT, and $d$ is the distance of the detector from the beam-dump.
 The values for the experimental parameters are summarized in table 5.
@@
 Target Mass Distance POT $\langle\pi/\rm{POT}\rangle$
-argon 24.4 kg 27.5 m $1.38\times 10^{23}$ 0.09
-cesium-iodide 14.6 kg 19.3 m $3.20\times 10^{23}$ 0.0848
+Ar 24.4 kg 27.5 m $1.38\times 10^{23}$ 0.09
+CsI 14.6 kg 19.3 m $3.20\times 10^{23}$ 0.0848
 The resulting expected counts for the two detectors’ runs are shown in Fig.
 4.
-The uncertainty in the number of counts is determined from two sources: the uncertainty in the cross section (calculated as above), and from the experimental systematic uncertainty (which is dominated by the 10% uncertainty in the total neutrino flux).
-We take our systematic uncertainty from COHERENT’s analysis (not including their theoretical uncertainties): 13% for argon and 12% for cesium iodide.
-Overall we achieved good agreement reproducing COHERENT’s results, with all predictions well within the errors for both COHERENT’s predictions and experimental results.
-However, without knowledge of their exact theoretical model it is difficult to reproduce the result precisely.
+The uncertainty in the number of counts is determined from two sources: the uncertainty in the cross section (calculated as above), and from the experimental systematic uncertainty (which is dominated by the 10% uncertainty in the total neutrino flux, but also includes detector effects such as efficiency cuts and quenching factors).
+We take our systematic uncertainty from COHERENT’s analyses (not including their theoretical uncertainties): 13% for argon and 12% for cesium iodide [43, 2].
+Overall we find good agreement with COHERENT’s results, with all predictions well within the errors for both COHERENT’s predictions and experimental results.
+However, without knowledge of their exact theoretical cross section, it is difficult to reproduce their result precisely.
 In this instance the result obtained using the form factor derived from the electric-charge distribution provides a useful benchmark to compare against.
@@
 As with the argon cross sections, the different form factors all produce very similar results because argon-40 is an even-even nucleus.
-It is worth noting that the reduction in the theoretical uncertainty achieved for our cross section predictions has essentially no effect on the total uncertainty of our event rate predictions, since they are dominated by the uncertainty in the neutrino flux.
-VI Conclusion
+It is worth noting that the reduction in the theoretical uncertainty achieved for our cross section predictions has essentially no effect on the total uncertainty of our event rate predictions, since they are dominated by the systematic uncertainty in the total neutrino flux.
+V Conclusion
 In this work we have computed precise flux-averaged cross section predictions for the CEvNS experiments being carried out by the COHERENT collaboration using argon, germanium and cesium-iodide targets.
 These predictions were made using both Skyme-Hartree-Fock calculations of weak radii, and shell model calculations of nuclear structure factors.
-The shell model results included sub-leading nuclear effects such as from spin-orbit coupling and axial currents.
-We used the Skyme-Hartree-Fock calculation to provide a cross-check of the density distribution obtained with the shell model.
-The results for the $F_{M}$ response showed very good agreement between these two methods.
-Our results are in broad agreement with prior work but have substantially reduced the theoretical errors by up to a factor of 10.
+These were compared with an empirically derived cross section prediction based on charge-radius measurements.
+We find that the our calculated weak radii and $F_{M}$ nuclear responses showed reasonable agreement between the two methods and the charge radii data.
+While the shell model results included additional nuclear effects such as from spin-orbit coupling and axial currents, they were not found to alter the predictions significantly.
+We observe a small difference between the the SHF and shell model calculations, where the latter gives a slightly smaller cross section for germanium and iodine targets.
+Given that spin-orbit and axial effects are negligible, we conclude that this difference is entirely driven by the density distribution.
+These results are in broad agreement with prior work, but have substantially smaller theoretical errors, even after accounting for the systematic difference between our SHF and shell model results.
+In the case of CsI, the difference between our SHF and shell model results is a factor of 2 times smaller than COHERENT’s prediction.
+Further work should be done to understand the origin of this discrepancy and implement more advanced uncertainty estimates.
 Future advances on the experimental side will both reduce the uncertainty on the neutrino flux and increase the number of detected CEvNS events, reducing statistical errors.
@@
 Appendix A Point-like nucleon distributions
-Full results for the point-like distribution of nucleons in each nuclear target are given in the tables below.
+Full results for the point-like distribution of nucleons in each nuclear target are given in tables 6 and 7.
 The SHF results were used to construct Fermi-form factors for the Weak-charge distribution.
@@
 with an equivalent form for the neutrons.
-Table 6: Skyrme-Hartree-Fock predictions of root mean square point like proton and neutron radii.
-Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm)
-40Ar SkM 3.356 3.445
-SLy6 3.330 3.412
-SV-min 3.321 3.418
-UNEDF1 3.296 3.407
-Average 3.326(25) 3.420(17)
-70Ge SkM 3.949 4.003
-SLy6 3.944 3.988
-SV-min 3.922 3.980
-UNEDF1 3.904 3.975
-Average 3.930(21) 3.986(12)
-72Ge SkM 3.965 4.053
-SLy6 3.959 4.031
-SV-min 3.940 4.031
-UNEDF1 3.922 4.028
-Average 3.946(20) 4.036(12)
-73Ge SkM 3.973 4.077
-SLy6 3.965 4.054
-SV-min 3.947 4.054
-UNEDF1 3.931 4.053
-Average 3.954(19) 4.059(11)
-74Ge SkM 3.980 4.099
-SLy6 3.976 4.081
-SV-min 3.955 4.077
-UNEDF1 3.941 4.078
-Average 3.963(18) 4.084(10)
-76Ge SkM 4.016 4.163
-SLy6 4.004 4.138
-SV-min 3.978 4.129
-UNEDF1 3.975 4.140
-Average 3.993(20) 4.143(14)
-127I SkM 4.702 4.836
-SLy6 4.698 4.824
-SV-min 4.676 4.814
-UNEDF1 4.676 4.822
-Average 4.688(14) 4.824(9)
-133Cs SkM 4.763 4.898
-SLy6 4.759 4.885
-SV-min 4.739 4.878
-UNEDF1 4.740 4.887
-Average 4.750(12) 4.887(8)
+Table 6: Skyrme-Hartree-Fock predictions of root mean square point like proton, neutron, and weak radii.
+Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm) $R_{w}$ (fm)
+40Ar SkM 3.356 3.445 3.539
+SLy6 3.330 3.412 3.507
+SV-min 3.321 3.418 3.513
+UNEDF1 3.296 3.407 3.503
+70Ge SkM 3.949 4.003 4.083
+SLy6 3.944 3.988 4.068
+SV-min 3.922 3.980 4.061
+UNEDF1 3.904 3.975 4.057
+72Ge SkM 3.965 4.053 4.134
+SLy6 3.959 4.031 4.112
+SV-min 3.940 4.031 4.113
+UNEDF1 3.922 4.028 4.111
+73Ge SkM 3.973 4.077 4.159
+SLy6 3.965 4.054 4.135
+SV-min 3.947 4.054 4.136
+UNEDF1 3.931 4.053 4.136
+74Ge SkM 3.980 4.099 4.181
+SLy6 3.976 4.081 4.163
+SV-min 3.955 4.077 4.160
+UNEDF1 3.941 4.078 4.161
+76Ge SkM 4.016 4.163 4.245
+SLy6 4.004 4.138 4.220
+SV-min 3.978 4.129 4.212
+UNEDF1 3.975 4.140 4.223
+127I SkM 4.702 4.836 4.907
+SLy6 4.698 4.824 4.895
+SV-min 4.676 4.814 4.886
+UNEDF1 4.676 4.822 4.894
+133Cs SkM 4.763 4.898 4.969
+SLy6 4.759 4.885 4.955
+SV-min 4.739 4.878 4.949
+UNEDF1 4.740 4.887 4.958
 Table 7: Shell model predictions of the point-proton, point-neutron and weak radii, using the various shell model calculations for each of the isotopes considered.
@@
 GCN5082* 4.72493 4.99011 5.06455
-Appendix B Comparison to nuclear observables
-Here we compare experimental nuclear observables to the theoretical shell model values, namely the nuclear energy levels, as well as the magnetic dipole and electric quadrupole moments and transitions.
-For each of these observables, we provide the values for 40Ar, 70,72,74,76Ge and 133Cs below, whilst the 73Ge and 127I comparisons can be found in [39].
-B.1 Argon-40
-Figure 5: Experimental energy levels (keV) for 40Ar compared with levels from the SDPF-NR, SDPF-U, SDPF-MU and EPQQM shell model calculations.
+Appendix B Details of the shell model calculations and comparison with experiment
+Here we provide further details of the shell model calculations, including valence spaces and potential truncations, and nuclear interactions.
+We also compare experimental nuclear observables to the theoretical shell model values, namely the nuclear energy levels, as well as the magnetic dipole and electric quadrupole moments and transitions.
+For each of these observables, we provide the values for 40Ar, 70,72,74,76Ge and 133Cs below, whilst the 73Ge and 127I comparisons can be found in reference [41].
+Information on the ground state is most important for comparison with theoretical values, as it is used in the scattering calculations here.
+However, as the ground state of even-even nuclei is $0^{+}$, the magnetic dipole and electric quadrupole moment values are identically zero.
+Nevertheless, we employ the next non-zero state ($2_{1}^{+}$), which can give an indication of the overall accuracy of the calculation through information on the spectrum beyond the ground state.
+40Ar
+40Ar shell model calculations were performed in reference [41] with the $sdpf$ valence space, which constitutes single particle levels $1d_{5/2}$, $2s_{1/2}$, $1d_{3/2}$, $1f_{7/2}$, $2p_{3/2}$, $1f_{5/2}$ and $2p_{1/2}$, and assuming a 16O core.
+The valence space truncation employed consisted of an unrestricted $sd$ shell and an empty $pf$ shell for the protons, and a full $sd$ shell with an unrestricted $pf$ shell for the neutrons.
+Four shell model interactions were used: SDPF-NR [49], SDPF-U [50], EPQQM [51] and SDPF-MU [52].
+The 40Ar theoretical energy levels are plotted against the experimental values in Fig.
+5.
+The first two theoretical energy levels for all interactions are consistent with experiment within $\sim 200$ keV, except for SDPF-MU which is within $\sim 400$ keV from experiment for the $2_{1}^{+}$ state.
+Overall, the energy spectrum of the EPQQM interaction is least consistent with the other shell model calculations and with experiment, with the spectrum being much more sparse and spread out in the level spacings.
+Note that all calculations miss the first-excited 0+ state, which is due to a collective multiparticle-multihole excitation outside the valence space and will be remarked upon further in the following comparison with electromagnetic observables.
+The magnetic dipole and electric quadrupole moments and transitions are presented in Table 8, where the effective $g$ factors and charges employed take on values of $e_{p}=1.5$, $e_{n}=0.5$, $g_{sp}=5.586$, $g_{sn}=-3.826$, $g_{lp}=1$, and $g_{ln}=0$.
+It is found that the shell model moments are not consistent with experimental values, and most experimental transitions are poorly reproduced as well.
+The reason for the failure of the shell model for the case of 40Ar is understood: the assumption that 40Ca is an inert core is not justified.
+Its ground state is near spherical, as assumed in the shell model, but it has low-excitation deformed states formed by exciting pairs of protons and neutrons across the shell gaps at $N=20$ and $Z=20$.
+This leads to mixing between the spherical shell model states and the deformed states in nuclei like 40Ar.
+The case of 42Ca, an isotone of 40Ar, is best studied, as has recently been reviewed by Stuchbery and Wood [53].
+They infer that the first 2+ state of 42Ca is about 50% due to the spherical shell model configuration.
+The ground state is about 65% of the shell model configuration.
+Similar mixing is present in 40Ar.
+Thus the shell model calculations performed here are limited by the presence of configuration mixing well beyond the usual valence space.
+Figure 5: Experimental energy levels (keV) for 40Ar are plotted up to 4000 keV and are compared with levels from the SDPF-NR, SDPF-U, SDPF-MU and EPQQM shell model calculations.
 Table 8: Magnetic dipole and electric quadrupole moments, and electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions, between low-lying states of 40Ar.
-Experimental transition values from [56], while moments taken from [57, 58].
+Experimental transition values are from [54], while moments are taken from [55, 56].
 Q [$e$fm2] $\mu$ [n.m.]
@@
 $2_{2}^{+}\rightarrow 2_{1}^{+}$ 23.76 17.83 28.13 60.55 146.26(4063) 0.2710 0.3055 0.2952 0.1203 0.06623(1074)
-B.2 Germanium-70,72,74,76
-Figure 6: Low-lying experimental energy levels (keV) for germanium isotopes compared with levels from shell model calculations using the JUN45 and jj44b interactions.
-Table 9: Magnetic dipole and electric quadrupole moments, for low-lying states of 70,72,74,76Ge. Experimental moments taken from [57, 58, 59].
+70,72,73,74,76Ge
+The germanium stable isotopes were studied in the unrestricted $f_{5}pg_{9}$ model space ($2p_{3/2}$, $1f_{5/2}$, $2p_{1/2}$ and $1g_{9/2}$ single-particle levels) with the JUN45 [57] and jj44b [58] shell model interactions in Ref.
+[41].
+In addition, the GCN2850 interaction [59] was used in reference [38, 42] with a valence space truncation that consisted of limiting the occupation number of the $1g_{9/2}$ level to no more than two nucleons above the minimum occupation for all isotopes.
+The experimental energy spectra are plotted against JUN45 and jj44b results for the even isotopes in Fig.
+6.
+In the cases of 70,72,74Ge, the low-lying experimental spectra are a little better reproduced by the JUN45 interaction, which, unlike jj44b, was fitted to the Ge isotopes in consideration.
+The shell model spectra for 76Ge are practically identical to one another up to $\sim 1.9$ MeV.
+An agreement between theoretical and experimental levels of $\sim 200$ keV would be considered satisfactory for such shell model calculations using heavier nuclei, however this agreement exists only for some levels of the even isotopes.
+See also [41] for 73Ge results, where the theoretical spectra were found to be inconsistent with another, and not particularly consistent with experiment.
+The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided for the even isotopes in Tables 9 and 10.
+For the magnetic dipole moments of the even isotopes, the experimental $2_{1}^{+}$ and $4_{1}^{+}$ states are well reproduced by both shell model interactions, whereas the jj44b value of the $2_{2}^{+}$ state is consistently closer to experiment compared to JUN45.
+Note that the compiler, Stone [55], has increased the uncetainties on the magnetic moments of the $2_{1}^{+}$ states reported by McCormick et al.
+[60] based on their statement that the absolute scale of the moments remains uncertain on the level of 20%.
+The $2_{1}^{+}$ electric quadrupole moment is quite inconsistent with the experimental values for 70,72,74Ge, where only the jj44b 76Ge value lies within the experimental range.
+Such discrepancies are not uncommon in medium-mass and heavy nuclei: the shell model is better equipped to describe single-nucleon features of the wavefunction than collective features like the shape of the charge distribution.
+Figure 6: Experimental energy levels (keV) for germanium isotopes are plotted up to 2500 keV (70Ge), 2100 keV (72Ge), 2165 keV (74Ge) and 2200 keV (76Ge), and compared with levels from shell model calculations using the JUN45 and jj44b interactions.
+Table 9: Magnetic dipole and electric quadrupole moments for low-lying states of 70,72,74,76Ge. Experimental moments are taken from [55, 56, 61, 60].
 70Ge 72Ge
@@
 $4_{1}^{+}$ +16.89 -10.15 +0.81 +0.91 +1.3(4) -0.76 -18.06 +1.20 +0.87 +0.8(6)
-Table 10: Electric quadrupole [B(E2)] transitions between low-lying states of 70,72,74,76Ge. Experimental transitions taken from [60, 61, 62, 63].
+Table 10: Electric quadrupole [B(E2)] transitions between low-lying states of 70,72,74,76Ge. Experimental transitions are taken from [62, 63, 64, 65].
 70Ge 72Ge 74Ge 76Ge
@@
 $4_{1}^{+}\rightarrow 2_{1}^{+}$ 772.7 885.4 756.6(554) 738.3 809.7 698.0(153)
-B.3 Cesium-133
-Figure 7: Experimental energy levels (keV) for 133Cs, compared with levels from SN100PN and GCN5082 shell model calculations for two different valence space truncations.
-Table 11: Magnetic dipole and electric quadrupole moments for low-lying states of 133Cs.
-Experimental moments taken from [64, 57, 58].
-GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
+127I
+Shell model calculations for 127I were performed in reference [41] in the model space which includes all proton and neutron orbits in the major shell between magic numbers 50 and 82, with single particle levels $1g_{7/2}$, $2d_{5/2}$, $2d_{3/2}$, $3s_{1/2}$ and $1h_{11/2}$, and with the SN100PN [66] and GCN5082 [59] interactions.
+The model space restrictions are described in reference [41], with the four different calculations given the names “SN100PN” and “GCN5082” (which share the same truncation), as well as “Truncated SN100PN” (with a more truncated scheme), and “GCN5082 Expanded” (with an alternative truncation to “GCN5082”).
+The valence space restriction employed in reference [38, 42] with the Baldridge and Dalton (B&D) interaction [67] involves fixing the occupation number of $1h_{11/2}$ to the minimum allowed nucleon number.
+See Ref.
+[41] for a comparison between shell model predictions and experimental nuclear structure data.
+133Cs
+Here, we consider the same valence space and shell model interactions (SN100PN and GCN5082) as 127I.
+We have performed four calculations for 133Cs, two of which employ the same truncations as the “SN100PN” and “GCN5082” calculations for 127I (and are given the same names here).
+In order to investigate the sensitivity of the shell model predictions to the truncation scheme, we consider two additional calculations, one each for the SN100PN and GCN5082 interactions, where the neutron $2d_{5/2}$ level now has a minimum of 4 neutrons (as opposed to previously being completely filled), with the remainder of the proton and neutrons levels retaining the previous truncations.
+This is to explore the effect of a less truncated valence space on the final results.
+These two additional calculations are termed “SN100PN Expanded” and “GCN5082 Expanded”.
+The experimental energy spectra are plotted against the theoretical values for all the above calculations in Fig.
+7.
+Only the SN100PN result reproduces the correct ground state spin and parity of $7/2^{+}$, however all four calculations reproduce the correct spins and parities of the three lowest-lying states within $200$ keV of experiment.
+For both shell model interactions, using a less truncated valence space causes the spacing of the levels to widen compared to the original calculations, generally pushing the spectrum further up.
+This is likely due to the opening up of the $2d_{5/2}$ level in these calculations, which increases pairing correlations in the ground state of the even-even configuration space that remains with the odd nucleon removed.
+The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided in Tables 11 and 12.
+The proton and neutron effective charges and $g$ factors employed here are the same as those used in reference [41] for 127I.
+The experimental magnetic dipole moments are much better reproduced by the shell model calculations compared to the electric quadrupole ones.
+The experimental electric quadrupole transitions are only reproduced within the uncertainty margins by the GCN5082 calculation, specifically for transitions from the $5/2_{1}^{+}$ and $3/2_{1}^{+}$ states to the ground state.
+On the other hand, the theoretical magnetic dipole transition values are not consistent with experiment.
+However, the important test of the shell model wavefunction is the magnetic dipole moment, which is sensitive to the wavefunction of the single state, and to the single nucleon structure of the state, rather than to emerging collective properties, which affect the E2 transition rates and the quadrupole moments.
+The magnetic dipole transitions are suppressed among the dominant low-excitation configurations, as is evident from their small values, and so become very sensitive to small components in the wavefunction that are not of interest here.
+Despite the lack of agreement for some observables, the present shell model calculations can be trusted to include the important single-particle configurations in the ground-state wavefunction.
+Figure 7: Experimental energy levels (keV) are plotted up to 750 keV for 133Cs, compared with levels from SN100PN and GCN5082 shell model calculations for two different valence space truncations (#2 indicates “Expanded” truncations).
+Table 11: Magnetic dipole and electric quadrupole moments for low-lying states of 133Cs. Experimental moments are taken from [68, 55, 56, 69]. “#2” refers to the expanded truncation schemes.
 Q [$e$fm2] $\mu$ [n.m.]
-State SN100PN SN100PN* GCN5082 GCN5082* Exp. SN100PN SN100PN* GCN5082 GCN5082* Exp.
+State SN100PN #2 GCN5082 #2 Exp. SN100PN #2 GCN5082 #2 Exp.
 $7/2_{\text{gs}}^{+}$14.11 -13.28 -18.71 -17.13 -0.343(10) 2.6649 2.6880 2.6715 2.7222 +2.5778(14)
-$5/2_{1}^{+}$66.40 -67.47 -63.36 -60.36 30(2) [^1] Sign not specified 3.54 3.62 3.36 3.50 +3.45(2)
+$5/2_{1}^{+}$66.40 -67.47 -63.36 -60.36 --22(5) [^1] From [69]. A different value, without sign, is given in Ref. [56]. 3.54 3.62 3.36 3.50 +3.45(2)
 $5/2_{2}^{+}$ 2.22 2.19 2.38 2.33 +2.25(15)
 Table 12: Electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions between low-lying states of 133Cs.
-Experimental transition values from [65].
-GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
+Experimental transition values are taken from [70].
+“#2” refers to the expanded truncation schemes.
 B(E2) [$e^{2}$fm4]
-Transition SN100PN SN100PN* GCN5082 GCN5082* Exp.
+Transition SN100PN #2 GCN5082 #2 Exp.
 $5/2_{1}^{+}\rightarrow 7/2_{\text{gs}}^{+}$ 178.7 155.2 242.8 177.2 233.9(161)
@@
 Appendix C EFT/Nuclear Operators
-The nuclear operators employed in Eq. (IV.1) have the form
+The nuclear operators employed in Eq. (III.1) have the form
 $\begin{split}M_{JM}(q\vec{x})&\equiv j_{J}(qx)Y_{JM}(\Omega_{x}),\\ \Sigma^{\prime}_{JM}(q\vec{x})&\equiv-i\left[\frac{\vec{\nabla}}{q}\times\vec{M}^{M}_{JJ}(q\vec{x})\right]\cdot\vec{\sigma}_{N}=[J]^{-1}\left[-\sqrt{J}\ \vec{M}^{M}_{JJ+1}(q\vec{x})+\sqrt{J+1}\ \vec{M}^{M}_{JJ-1}(q\vec{x})\right]\cdot\vec{\sigma}_{N},\\ \Phi^{\prime\prime}_{JM}(q\vec{x})&\equiv i\left(\frac{\vec{\nabla}}{q}M_{JM}(q\vec{x})\right)\cdot\left(\vec{\sigma}_{N}\times\frac{1}{q}\vec{\nabla}\right),\end{split}$ (34)
@@
 $\begin{split}&\langle j^{\prime}m^{\prime}|T_{JM}|jm\rangle=(-1)^{j^{\prime}-m^{\prime}}\begin{pmatrix}\vspace{2mm}j^{\prime}&J&j\\ \vspace{2mm}-m^{\prime}&M&m\hskip 5.69054pt\end{pmatrix}\langle j^{\prime}||T_{J}||j\rangle.\end{split}$ (35)
-The one-nucleon matrix elements in Eq. (IV.1) have analytic expressions of the form [15]:
+The one-nucleon matrix elements in Eq. (III.1) have analytic expressions of the form [18]:
 $\Big{\langle}n^{\prime}l^{\prime}\frac{1}{2}j^{\prime}\Big{|}\Big{|}M_{J}\Big{|}\Big{|}nl\frac{1}{2}j\Big{\rangle}$ $=\langle n^{\prime}l^{\prime}|j_{J}(qr_{i})|nl\rangle(-1)^{j+1/2+J}\sqrt{\frac{1}{4\pi}}\,\bigl{[}(2j^{\prime}+1)(2j+1)\bigr{]}^{\frac{1}{2}}\bigl{[}(2J+1)(2l+1)(2l^{\prime}+1)\bigr{]}^{\frac{1}{2}}$
@@
 Appendix D Form factors
-In this section we plot the form factors for the individual responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ using the shell model calculations performed in this work, where the $\Sigma^{\prime}$ response is only plotted for isotopes with non-zero nuclear ground state spin.
+In this section we plot the form factors for the individual responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ using the shell model results, where the $\Sigma^{\prime}$ response is only plotted for isotopes with non-zero nuclear ground state spin.
 This is presented for 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs, highlighting the corresponding COHERENT range of interest for each of these nuclei.
-The form factor expressions employed to plot these are presented in [39], except for 133Cs which is provided in Appendix D.
+The form factor expressions employed to plot these are presented in reference [41], except for 133Cs which is provided in Appendix D.
 The KN and SHF form factors were plotted using $(ZF\left(q,r_{\text{rms}}^{p})+NF(q,r_{\text{rms}}^{n})\right)^{2}$, while the shell model counterparts were plotted using $\sum_{N,N^{\prime}=p,n}F^{(N,N^{\prime})}_{X,X}(q^{2})$, where $F^{(N,N^{\prime})}_{X,X}(q^{2})$ is defined in Eq.
@@
 Figure 9: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.2 Germanium-70
+The maximum difference between the form factors for the $M$ response at the minimum $q$ value is $\sim 0.3\%$, and $\sim 11\%$ at the maximum $q$ value.
+For $\Phi^{\prime\prime}$, these values are $\sim 24\%$ and $\sim 21\%$, respectively, where the largest differences occur between the SDPF-MU and EPQQM interactions.
+D.2 Germanium isotopes
+The $M$ response is consistent amongst all calculations, with a maximum difference of $\sim 0.2\%$ at the lower end of the COHERENT $q$ region-of-interest (ROI) for the even isotopes, and $\sim 3\%$ at the higher end for 70,72Ge.
+For 73Ge these take on the values of $\sim 0.3\%$, and $\sim 20\%$ at the higher end for 73,74,76Ge.
+For the $\Phi^{\prime\prime}$ channel, the maximum difference is observed between the GCN2850 and jj44b shell model calculations, which is $\sim 80\%$ for 70Ge at both ends of the $q$ spectrum, factor of $\sim 2.4$ for 72Ge, factor of $\sim 2.3$ for 73Ge, factor of $\sim 2.2$ for 74Ge, and $\sim 60\%$ for 76Ge.
+The difference in the SD $\Sigma^{\prime}$ channel for 73Ge is $\sim 30\%$ at both ends of the $q$ ROI.
 Figure 10: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 11: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.3 Germanium-72
 Figure 12: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 13: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.4 Germanium-73
 Figure 14: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 16: $\Sigma^{\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.5 Germanium-74
 Figure 17: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 18: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.6 Germanium-76
 Figure 19: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 20: $\Phi^{\prime\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.7 Iodine-127
+D.3 Iodine-127
+The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 30\%$ at the maximum $q$ value.
+For $\Phi^{\prime\prime}$, these values are $\sim 55\%$ and $\sim 50\%$, respectively.
+For $\Sigma^{\prime}$, the factor differences are $\sim 44$ and $\sim 81$, due to the large difference between the Truncated SN100PN calculation and the remaining calculations.
+This is consistent with the integrated form factor (IFF) results in reference [41], which showed a more pronounced difference between the SD Truncated SN100PN IFF results and other calculations.
+This indicates that a more severe valence space truncation for a shell model interaction has a larger effect compared to a second more comparable truncation.
 Figure 21: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 Figure 23: $\Sigma^{\prime}$ form factors employing the different shell model interactions used for this isotope, plotted against the COHERENT range-of-interest.
-D.8 Cesium-133
+D.4 Cesium-133
+The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 20\%$ at the maximum $q$ value.
+For $\Phi^{\prime\prime}$, these values are $\sim 6\%$ and $\sim 7\%$, respectively.
+For $\Sigma^{\prime}$, these differences are $\sim 13\%$ and $\sim 18\%$.
 Figure 24: SI ($M$) form factors employing the different shell model interactions used for this isotope.
@@
 133Cs Form Factor Expressions
-The explicit nuclear form factors for 133Cs are given below using the convention of [35]:
+The explicit nuclear form factors for 133Cs are given below using the convention of Fitzpatrick et al. [38]:
 $F^{(N,N^{\prime})}_{X,Y}(q^{2})\equiv\frac{4\pi}{2J_{i}+1}\sum_{J=0}^{2J_{i}}\langle J_{i}||X_{J}^{(N)}||J_{i}\rangle\langle J_{i}||Y^{(N^{\prime})}_{J}||J_{i}\rangle,$ (41)
-and can be used to obtain the $\mathcal{F}_{N}^{Y}(q^{2})$ in Eq.
-(IV.1).
-Since no interference occurs between operators in CEvNS, we simplify the notation and write $F^{(N,N^{\prime})}_{X,X}(q^{2})\equiv F^{(N,N^{\prime})}_{X}(q^{2})$.
+and can be used to obtain the $\mathcal{F}_{N}^{\Sigma^{\prime}}(q^{2})$ responses in Eq. (III.1) through the relation:
+$|\mathcal{F}^{\Sigma^{\prime}}_{N}|^{2}(q^{2})=\frac{1}{4}\frac{2J_{i}+1}{4\pi}F_{\Sigma^{\prime}}^{N,N}$ (42)
+This can be used to show that, in the isospin basis, the spin structure functions are:
+$S^{\mathcal{T}}_{00}$ $=$ $\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{00}(\mathbf{q}^{2})$
+$S^{\mathcal{T}}_{11}$ $=$ $(1+\delta^{\prime}(\mathbf{q}^{2}))^{2}\left(\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{11}(\mathbf{q}^{2})\right)$
+$S^{\mathcal{T}}_{01}$ $=$ $2(1+\delta^{\prime}(\mathbf{q}^{2}))\left(\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{01}(q^{2})\right)$
+While describing CEvNS only requires the operators $M,\Phi^{\prime\prime}$ and $\Sigma^{\prime}$, below we provide the form factors for the full list of operators in [38].
 SN100PN
@@
 $F_{\Phi^{\prime\prime}}^{(p,p)}=e^{-2y}(91.3-223.y+207.y^{2}-92.9y^{3}+21.6y^{4}-2.55y^{5}+0.131y^{6}-0.00154y^{7}+5.26\times 10^{-6}y^{8})$
-$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.015-0.00813y+0.000808y^{2}+0.00263y^{3}-(8.21*10^{(}-6))y^{4}-0.0000477y^{5}+0.0000242y^{6}+5.25\times 10^{-7}y^{7}+4.03\times 10^{-9}y^{8})$
+$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.015-0.00813y+0.000808y^{2}+0.00263y^{3}-8.21\times 10^{-6}y^{4}-0.0000477y^{5}+0.0000242y^{6}+5.25\times 10^{-7}y^{7}+4.03\times 10^{-9}y^{8})$
 $F_{\Delta}^{(p,p)}=e^{-2y}(2.74-6.58y+6.44y^{2}-2.97y^{3}+0.703y^{4}-0.0812y^{5}+0.00379y^{6}-2.48\times 10^{-6}y^{7}+5.93\times 10^{-10}y^{8})$
@@
 $F_{M\Phi^{\prime\prime}}^{(n,p)}=e^{-2y}(-743.+2510.y-3250.y^{2}+2110.y^{3}-751.y^{4}+150.y^{5}-16.3y^{6}+0.848y^{7}-0.015y^{8}+0.0000709y^{9})$
-$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.00497+0.0599y-0.158y^{2}+0.164y^{3}-0.0838y^{4}+0.0234y^{5}-0.00359y^{6}+0.00028y^{7}-8.51\times 10^{-6}y^{8}+2.73\times 10^{-9}y^{9}$
+$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.00497+0.0599y-0.158y^{2}+0.164y^{3}-0.0838y^{4}+0.0234y^{5}-0.00359y^{6}+0.00028y^{7}-8.51\times 10^{-6}y^{8}+2.73\times 10^{-9}y^{9})$
 $F_{M}^{(n,n)}=e^{-2y}(6050.-26100.y+44400.y^{2}-39100.y^{3}+19600.y^{4}-5850.y^{5}+1040.y^{6}-105.y^{7}+5.58y^{8}-0.125y^{9}+0.000963y^{10})$
@@
 “Max Occ” refers to the maximum occupation number allowed for each orbital.
-GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
-SN100PN SN100PN* GCN5082 GCN5082*
+“#2” refers the “Expanded” truncation schemes.
+SN100PN SN100PN(#2) GCN5082 GCN5082(#2)
 $n$ $l$ $j$ Max Occ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$ $n^{p}_{nlj}$ $n^{n}_{nlj}$